### v1

# Learning to Infer Parameterized Representations of Plants from 3D Scans

Abstract
Reconstructing faithfully the 3D architecture of plants from unstructured observations is a challenging task. Plants frequently contain numerous organs, organized in branching systems in more or less complex spatial networks, leading to specific computational issues due to self-occlusion or spatial proximity between organs. Existing works either consider inverse modeling where the aim is to recover the procedural rules that allow to simulate virtual plants, or focus on specific tasks such as segmentation or skeletonization. We propose a unified approach that, given a 3D scan of a plant, allows to infer a parameterized representation of the plant. This representation describes the plant’s branching structure, contains parametric information for each plant organ, and can therefore be used directly in a variety of tasks. In this data-driven approach, we train a recursive neural network with virtual plants generated using an L-systems-based procedural model. After training, the network allows to infer a parametric tree-like representation based on an input 3D point cloud. Our method is applicable to any plant that can be represented as binary axial tree. We evaluate our approach on Chenopodium Album plants, using experiments on synthetic plants to show that our unified framework allows for different tasks including reconstruction, segmentation and skeletonization, while achieving results on-par with state-of-the-art for each task.
Figure 1: Our method takes a 3D point cloud of a plant as input, and outputs a parameterized representation of the plant. This representation encodes the plant’s branching structure and geometry along with semantic information such as organ type, and allows for multiple tasks including reconstruction, organ segmentation and skeleton extraction.
11footnotetext: Inria centre at the University Grenoble Alpes22footnotetext: Inria centre of Lyon
1 Introduction
Inferring high-level information of plants from observations in a way that respects the biological traits (e.g. the radius of a lateral branch is never larger than that of it’s parent stem) is an important problem that has been studied for decades [52]. High-level information includes the reconstruction of the 3D geometry of a plant, the segmentation of a plant into parts, and the extraction of a skeleton or measurements. Observations are typically 2D images or 3D scans. Automatically extracting this information allows to use reconstructed 3D plants in content generation and virtual reality [29], and to use part segmentation, skeleton information, and measurements in plant phenotyping, e.g. [33, 13, 18].
The problem of inferring high-level information is challenging because plants have complex structures due to their branching systems, which leads to strong self-occlusions and ambiguities. Existing works that analyze 2D or 3D plant observations therefore either focus on inverse modeling, where the goal is to find growth rules that allow to generate a given plant, or are task-specific. Inverse modeling is challenging and existing methods are limited to simple branching structures [24, 48]. Task-specific methods focus on 3D reconstruction e.g. [29], organ segmentation e.g. [33], skeletonization e.g. [13], or measurement extraction e.g. [18].
We propose a unified approach that, given a 3D scan of a plant, allows to infer a parameterized representation of the plant. This representation contains parametric information on each plant organ, and can therefore be directly used to reconstruct the 3D geometry of the plant with corresponding segmentation labels and measurements, thereby allowing to solve a variety of tasks. We achieve this by learning a shape space of 3D plants based on representations generated by a biologically inspired procedural model, in a way that for any point in shape space, we can retrieve its corresponding parameterized representation. By learning a function that maps an observed 3D point cloud to the learned shape space, our work allows to find the representation of a plant without reconstructing the full procedural model rules. Hence, our work can be considered a step towards inverse modeling-based reconstruction.
The key idea that allows our method to address the challenging problem of inferring high-level information from 3D scan data plagued by occlusions is to learn a shape space with the help of instances generated using a procedural model made of recursive rules that describe the development of the plant architecture. We leverage a Lindenmayer-systems (L-systems) model [40] allowing to generate plants, and the recursive nature of plant growth in a recursive neural network to learn a latent distribution of the plants’ geometry and branching structure. We generate a synthetic dataset using procedural modeling rules, which allows for supervised learning from large amounts of occlusion-free training data. The synthetic plants are represented in a binary axial tree form that can be geometrically interpreted to obtain the 3D virtual plant. The binary tree format is exploited using a recursive auto-encoder. To address the problem of generalizing to 3D scans plagued by acquisition noise and occlusions, we virtually simulate the acquisition noise on the 3D virtual plants, in addition to simulated monocular depth maps during training and learn a mapping from simulated 3D scan data to the shape space.
We evaluate our method on synthetic Chenopodium Album plants, an annual plant, in early growth stages. We choose this plant because of the availability of real scans in dataset [34], and its simple structure in early stage that can be encoded into a binary tree representation. We use our method to perform the different tasks of 3D reconstruction, skeleton extraction and plant part segmentation. Our experiments show results that are on-par with the state-of-the-art, with more robustness to noisy and partial inputs. We also show qualitative results for instance segmentation, and discuss the potentials to apply our method on real scans.
The main contributions of this work are:
  • A unified approach to infer a parameterized representation of a plant from a 3D scan based on a recursive neural network.
  • 3D plant reconstruction, skeletonization, organ segmentation, with performance that is on-par with the state-of-the-art on each task.
  • A dataset of virtual plants, in a form of L-Strings, that represents instances of Chenopodium Album plants in early growth stages.
2 Related Work
Vision based systems and platforms have been developed for plant phenotyping and crops analysis [7, 8, 11, 10, 37]. Furthermore, plant scan datasets and data generation methods were provided for training data-driven models, e.g.  the rosebush annotated dataset ROSE-X [15], an annotated dataset of 3 plant species PLANEST-3D [31], and a labeled point cloud generation method of virtual plants [4]. Furthermore, small datasets of 3D scans were captured e.g. [34].
Existing works on plant phenotyping that deal with plant scan data belong to two main categories: works that aim for inverse plant modeling, and task-specific methods that aim to extract information. Our work is in between these categories, since it is centered around the inference of parameterized plant representations, which allows to extract various high-level information including the full branching structure and 3D geometry of the plant. This can be seen as a step towards inverse plant modeling as we infer an instance of a procedural model.
2.1 Inverse Plant Modeling
Procedural plant models generate a plant following growth rules. On the contrary, inverse plant modeling aims to find the rules that allow to generate the given plant. In other words, the goal is to infer the model given the plant geometry as input. Guo et al. [23] and Štáva et al. [47] infer procedural modeling rules generated using an L-system from 2D images of simple branching structures. Štáva et al. [48] introduce a parametric procedural model and estimate the parameters that allow this model to generate trees of a single species similar to an input 3D mesh. More recently, Lee et al. [24] used transformers to learn a latent space that allows to predict L-system rules that correspond to 3D point clouds of a tree, and to generate similar trees.
As Lee et al. [24], we use data-driven neural networks to learn a latent representation of the plant. However, we aim to directly predict the L-String that represents the input plant with its branching structure, geometrical and other biological traits, instead of predicting the L-system rules.
2.2 Task Specific Methods
Existing methods used in plant phenotyping from 3D input point clouds mainly focus on four tasks: 3D reconstruction of the plant, extracting a skeleton of the branching structure of the plant, segmenting the plant into organs, and extracting measurements from the point cloud.
3D Reconstruction
Reconstructing plant geometry from an input scan represented by a 3D point cloud is an important problem due to its applications in computer graphics and plant phenotyping. This is a particularly challenging 3D reconstruction problem as plants contain thin structures and as different organs lead to significant self-occlusions.
In computer graphics, Gonzalez et al. [22] use UnrealEngine [41] to reconstruct digital urban trees by reconstructing a mesh that represents the tree trunk, estimating the volume and density of the canopy, and filling the canopy with generated leaves. This method focuses on global geometry and ignores the branching structure of the plant architecture.
Another line of work focuses on the tree branching structure [14, 39, 54, 29, 53], by estimating a skeletal structure of the tree. Given the skeleton, a mesh surface can be reconstructed over each branch. However, the foliage part of the trees is either excluded from the input, or reconstructed by randomly generating leaves to fill the canopy.
Closest to our work, Prasad et al. [38] explore different implicit surface reconstruction algorithms, including classical and neural based ones, on 3D point clouds of a single plant. They demonstrate that this task is challenging due to fine geometric structures. Note that our method’s output contains more information than a 3D reconstruction, as our parametric output allows to infer information on the plant’s skeleton and organs. For 3D reconstruction, we compare our method experimentally to this work.
Skeletonization
Extracting the skeletal structure of plants is a well-studied problem. One line of works, discussed above, extracts skeletons as one part of a pipeline to reconstruct the plant [14, 39, 54, 29, 53].
Another line of works corresponds to optimization-based methods. Meyer et al. [32] use a point contraction algorithm to extract skeletons of leafless cherry trees. Works using graph-based approaches have shown to extract skeletal structures even in the presence of noise for trees [53, 30]. A recent deep learning-based approach uses a sparse convolutional neural network to estimate the medial axis of a tree [13]. Given a point cloud of a plant with an initial extracted skeleton, Chaudhury et al. [6] propose a stochastic optimization to refine the initial skeleton by estimating point transformations.
All discussed methods require a leafless input point cloud to estimate the skeleton. When leaves are present in the input point cloud, these methods estimate skeletal structures inside of the leaves, thereby adding noise to the output. Since our method outputs a segmented parameterized representation of the plant, extracting the skeleton of a leafy input plant does not pose additional challenges. We experimentally compare our method to one successful approach from each category for which code is available, namely [53, 30, 6].
Segmentation
Segmenting different organs from point clouds of plants is critical for phenotyping on both semantic and instance levels. Both learning and optimization-based techniques were proposed in the literature to segment a plant into organs. Mirande et al. [33] propose a graph-based optimization approach with botanical knowledge refinement for semantic and instance segmentation. Wahabzada et al. [51] present an unsupervised data-driven method using histogram clustering in a Euclidean space. For supervised learning methods, a benchmark for plant organ segmentation [49] has been released based on the ROSE-X dataset [15]. More complex deep learning architectures designed for organ segmentation lead to accurate results [26, 50, 25]. We compare our method experimentally to two of these successful approaches [26, 25].
Measurement Extraction
Extracting phenotypic measures like leaf area index or main stem height provides insight about plant and crop behavior in different conditions. Paproki et al. [36] propose a mesh processing technique for 3D plants phenotypic traits analysis from multi-view images. Chattopadhyay et al. [3] model and reconstruct apple trees with diameter estimation for the application of automatic dormant pruning. Ghahremani et al. [18] apply the RANSAC algorithm to model plant organs that fit an input point cloud to estimate phenotypic features like angles and diameters of leaves and branches.
Positioning
Our experiments focus on the three tasks of 3D reconstruction, skeletonization and segmentation. As shown in Table 1, only few works perform the three tasks of 3D reconstruction, skeletonization and segmentation, like [29, 54]. However, these methods are designed to be applied on big trees with trunks and foliage, and are not applicable on small annual plants where the leaf size to stem diameter proportion is different. On annual plants, our method remains generic due to the learned parametric model that contains the information needed for different tasks.
Work Recon- Skeletoni- Segment- Annual
struction zation ation plants
Gonzalez
 et al. [22]
Liu
 et al. [29]
Du
 et al. [14]
Linvy
 et al. [30]
Preuksakarn
 et al. [39]
Parsad
 et al. [38]
Dobbs
 et al. [13]
Chaudhury
 et al. [6]
Yan
 et al. [54]
Meyers
 et al. [32]
Mirande
 et al. [33]
Turgut
 et al. [49]
Turgut
 et al. [50]
Wahabzada
 et al. [51]
Li
 et al. [26]
Li
 et al. [25]
Paproki
 et al. [36]
Ghahremani
 et al. [18]
Ours
Table 1: Positioning of our method w.r.t. the ability to perform 3D reconstruction, skeletonization and part segmentation, and the applicability on small annual plants.
3 Background
We propose a method that learns a shape space of a plant species with the help of recursive neural networks. To learn this shape space, we rely on a large dataset of 3D plants with corresponding parametric representations that has been simulated with the help of plant models. This section provides background on plant modeling and on recursive neural networks.
3.1 Plant Modeling
Plant modeling aims to find a mathematical model that describes the complex geometry and growth rules for a species of plants. Moreover, plant models respect biological rules to allow realistic simulation of the plant’s appearance and behavior in different environmental conditions [5, 19]. The importance of plant modeling comes from its applications in biological and agronomical studies. The synthetic models can be used to perform growth simulation experiments, or estimate biological traits and growth patterns of plants. This allows to study plants with fast and non-invasive methods.
Procedural methods are among the most commonly used plant modeling approaches, including L-systems [40], the space colonization algorithm [43], and commercial tools such as SpeedTree. In our work, we use procedural plant models to generate data for training and evaluating our neural network.
Our implementation uses an L-system-based procedural approach [40] to generate plant data, a classical rule based plant modeling technique, that is particularly convenient due to the availability of tools such as L-Py [2]. In L-systems, a plant is represented as a string of symbols, called L-String, possibly bearing geometrical or biological parameters. Similarly to natural or computer language grammars, an L-system consists of a set of rewriting rules that define how plant components represented by L-String symbols change as time proceeds, by specifying how symbols get replaced by combinations of other symbols [40, 21]. At each step, the rules replace the symbols in parallel, resulting in a new L-String representing the next plant state.
L-Strings of plants mainly consist of two types of symbols: modules and brackets. Modules can refer to various plant parts e.g. stems, leaves, flowers, etc. Opening and closing brackets indicate the start and the end of every branch in the plant. Two successive modules in an L-sting have a parent-child relationship. Open brackets allow for a module to have more than one child, and for different children to have siblings relationships. However, each parent has at most one special child that corresponds to its successor on the same plant axis. Such L-Strings encode axial trees [40] that represent the plant’s architecture. Modules in L-Strings can have parameters that give information about a plant organ. This includes the organ’s age, order in the branch or geometric attributes (e.g.  length, radius, angle, etc.). By applying geometric interpretation rules on an L-String using turtle geometry [1], a 3D scene of the plant can be generated. The virtual plant comprises of connected 3D shapes corresponding to the modules. In other words, an L-String can encode a plant’s geometry in addition to its architecture.
Our L-system models are implemented using the L-Py [2] language, that provides an extension of the Python programming language to design procedural models using the L-systems formalism. L-Py allows to create and manipulate L-system rules, handle the information present in an L-String, and generate 3D virtual plants from L-Strings.
3.2 Recursive Neural Networks
Recursive Neural Networks (RvNN) [12] are a class of deep neural networks that apply the same network architecture recursively on input that is structured as a tree graph to give a single representation in a latent space. The term recursive refers to the network architecture being applied to the output of step $i-1$ during step $i$. The recursivity of RvNNs allows to handle input of varying size, as the network is applied in a bottom-up fashion from the leaves to the root. These networks learn hierarchical dependencies and correlations between input components.
Socher et al. [46, 45] introduced an RvNN that learns information on different domains, including natural language processing and natural scene image processing, by recursively parsing a tree graph with binary structure. Models based on RvNN have been applied to natural language processing [9], 3D indoor scene generation [28, 17], blood vessel synthesis [16], and 3D shape structure recovery [35], segmentation [55] and generation [27].
Of particular interest for our work are recursive auto-encoders for binary trees, such as the ones used to generate 3D shapes of man-made objects [27]. A recursive auto-encoder learns on a binary tree structure, where all nodes can be represented in a latent space of dimension $dim_{\mathcal{S}}$. The encoder follows the tree structure and recursively merges pairs of inputs to form a new point in the same latent space until the full tree is represented as a single point in latent space. Inversely, the decoder recursively decodes a single point into two points in the same latent space until the full tree structure is decoded.
Figure 2 shows an example of the network architecture of a recursive encoder-decoder pair. The encoder-decoder pair allows to learn correlations between continuous properties encoded in features along the tree. An RvNN typically contains different types of encoder-decoder pairs. This allows to implement different operations, such as merging or splitting different parts.
During decoding, given a point in latent space, an algorithm is needed to decide which decoder to apply and when to stop the recursion. To achieve this, the encoders and decoders are trained with ground truth structural information. During encoding, the information of which encoder to apply is available. For decoding, the RvNN jointly learns a set of classifiers that is applied before each decoding step.
Figure 2: The network architectures used for the recursive encoder and decoder. A binary tree structure is recursively encoded into latent space $\mathcal{S}$. Latent points $s_{1}\in dim_{\mathcal{S}}$ and $s_{2}\in dim_{\mathcal{S}}$ are merged into $s_{1,2}\in dim_{\mathcal{S}}$ by the encoder. Symmetrically, the decoder splits $s_{1,2}$ into two latent points $s_{1}$ and $s_{2}$. $dim_{h}$ denotes the dimension of the hidden layer.
L-Strings have tree-like hierarchical structure, and we use RvNNs to encode them into latent points while capturing the relationships between the different modules, and their features’ correlations along the tree. We also use RvNNs in the opposite direction, to retrieve the tree structure and its parameters.
4 Method
4.1 Overview
We aim to obtain a tree-like parametric representation $l$ in the form of an L-String from an input point cloud $P$ that stems from an unstructured 3D scan of a plant. Learning a direct regression from the space of 3D point clouds to the space $\mathcal{L}$ of L-Strings is a difficult problem, as the point clouds have varying number of points and are unstructured, and as the L-Strings varies both in discrete (i.e. number and type of modules) and continuous (i.e. values of the parameters describing angles and lengths) ways.
To address this problem, we combine an RvNN learned on $\mathcal{L}$ with an encoder that maps a point cloud to $\mathcal{S}$, as shown in Fig. 3. During training, we first learn the latent space $\mathcal{S}$ that allows to represent L-Strings of a fixed plant species. This allows to represent 3D plants of varying shapes, with different number of organs, and different skeletal structures using latent representations $s\in\mathcal{S}$, and is explained in Sec. 4.2. Second, we train a neural network called point cloud encoder that maps an input point cloud $P$ to a point $s\in\mathcal{S}$, as outlined in Sec. 4.3). During inference, the input point cloud $P$ is encoded in the latent space as a point $s$ using the point cloud encoder, and subsequently decoded to an L-String using the recursive decoder, as detailed in Sec. 4.4.
Figure 3: Overview: our method learns a latent space $\mathcal{S}$, that allows the mapping of 3D point clouds to L-Strings. At inference, the point cloud is mapped to $\mathcal{S}$ using the point cloud encoder on the left, and the resulting latent point allows to reconstruct the corresponding L-String using the L-String decoder on the right.
4.2 Representing L-Strings in Latent Space
We aim to learn a latent space $\mathcal{S}$ that allows to represent various static instances of the same species of plant. Plants are represented in the space of L-Strings $\mathcal{L}$ and can be generated with a set of L-System rules. However, the branching structure of the L-Strings may not always correspond to a binary tree structure. We therefore first simplify the L-Strings by summarizing the information of modules that always occur together in the plant species. The resulting combined modules are called nodes in the following, and we design the combination rules to guarantee a binary tree structure after combination. To simplify notations, we call the simplified L-String $l\in\mathcal{L}$ in the following.
Our goal is to learn an encoder function $E:\mathcal{L}\to\mathcal{S}$, and a decoder function $D:\mathcal{S}\to\mathcal{L}$, such that $l\approx D(E(l)),\;\;\forall l\in\mathcal{L}$.
Learning to encode different shape and structural information in latent points of fixed dimension is challenging. Inspired by [27], we use an RvNN to learn the hierarchical relations between the modules of the L-String, leveraging the recursive nature of plant structures [20].
To achieve this, each node of the tree is represented individually in a point in the latent space, $s_{i}\in\mathcal{S}$, using a node auto-encoder. In a second step, each subtree of the L-String is represented in $\mathcal{S}$ using a recursive auto-encoder. This procedure is carried out recursively from the tree-graph leaves up to its root.
4.2.1 Node Auto-encoders
Each type of node has a different set of parameters, e.g. angles, widths, radii. This results in nodes that have different dimensionality in general, and that are not directly comparable. To allow nodes with different number of parameters to be used as input in an RvNN, we first map the information of each node to $\mathcal{S}$. To achieve this, a node encoder-decoder pair is learned for each node type. In the following, let $E_{node,i}$ and $D_{node,i}$ denote the encoder and decoder for each type of node, with $i=1,\ldots,N$ and $N$ the number of types of nodes. Both $E_{node,i}$ and $D_{node,i}$ consist of a fully connected linear layer, followed by a $tanh$ activation.
4.2.2 Recursive Auto-encoders
Figure 4: The two criteria to merge/split points in latent space shown on an example tree. Merging is applied recursively in a bottom-up manner until the whole tree is merged into one point, while the splitting performs the inverse operation.
After applying node encoding, an input L-String is represented as a binary tree where all nodes are individually represented as points in $\mathcal{S}$. It can then serve as input to an RvNN auto-encoder. To recursively merge or split the nodes, we design encoder-decoder pairs based on the relationship of the input nodes in the tree. Two nodes to be merged can be either siblings or have a parent-child relationship in the tree.
Based on this observation, we consider the two merging/splitting criteria shown in Figure 4. For each criterion, we learn one recursive encoder-decoder pair. Both the sibling encoder-decoder $(E_{sib},D_{sib})$ and the parent-child encoder-decoder $(E_{pc},D_{pc})$ are implemented as shown in Fig 2.
4.2.3 Auxiliary Classifiers
In addition to the encoder-decoder pairs, it must be decided how to properly split the points to reconstruct the L-String structure. Each point in $\mathcal{S}$ is either a result of a merging process by one of the recursive encoders ($E_{sib}$, $E_{pc}$), or it represents an individual node. For the RvNN to decode structural information, it needs to learn to choose the appropriate decoder ($D_{sib}$, $D_{pc}$), or to end the recursive splitting. To predict the right decoder, we associate classes for the different splitting options (siblings, parent-child, stop) and jointly train a classifier on them.
Due to different node types, the network should needs to learn which node decoder $E_{node,i}$ to apply to reconstruct each node of the L-string. For that, a second classifier is trained, that allows to predict the node type for a latent point $s$ that is fully split. This classifier, called $C_{node}$ has $N$ output classes.
Both classifiers share the architecture of a two-layer fully connected neural network. The first layer maps the input latent point to a hidden representation using a linear transformation followed by a $tanh$ activation. The second layer maps the hidden features to an output vector that corresponds to the different classes.
4.2.4 Training
All auto-encoders are trained by minimizing a reconstruction loss at the node level. For node $n$ in L-String $l$, with a set of parameters, the reconstruction loss $L_{rec}(n)$ is the mean squared error between the input node parameter vector that contains the plant part information and its reconstruction. The reconstruction loss for the full L-String is
$L_{rec}=\sum_{n}L_{rec}(n).$ (1)
At each recursive encoding step $i$, the output point $s$ is a result of merging two points ($s_{1}^{(i)}$, $s_{2}^{(i)}$). Then in the inverse decoding step, $s$ will be split into ($\hat{s}_{1}^{(i)},\hat{s}_{2}^{(i)}$) that are supposed to represent the same information as ($s_{1}^{(i)}$, $s_{2}^{(i)}$) and should fall in the same region in $\mathcal{S}$. To learn a structured representation space into which we can map information from another modality (point clouds), we encourage corresponding latent points during encoding and decoding to be located close-by in $\mathcal{S}$ using a step-based recursion loss $L_{step}$ that is defined as
$L_{step}=\sum_{i}\left(\left(s_{1}^{(i)}-\hat{s}_{1}^{(i)}\right)^{2}+\left(s_{2}^{(i)}-\hat{s}_{2}^{(i)}\right)^{2}\right).$ (2)
The classifiers are trained with softmax classification and cross entropy loss. $C_{split}$ takes a latent point $s$ as input, and predicts one of the three classes: parent-child split, siblings split or leaf node (no split). $C_{node}$ predicts one of $N$ possible node types. We denote the cross entropy losses used to train these classifiers by $L_{split}$ and $L_{node}$, respectively.
Finally, the total loss that is used for training the recursive L-String auto-encoder is
$L_{total}=L_{rec}+L_{step}+L_{split}+L_{node}.$ (3)
4.3 Point Cloud Encoder
After representing L-Strings in latent space $\mathcal{S}$ has been learned, 3D plants can be represented as latent points $s\in\mathcal{S}$, and $s$ can be decoded into an L-String parametric representation $l$. To allow retrieving an L-String $l$ from an input point cloud $P$, we learn a new mapping function $E_{points}$ from the space of point clouds to $\mathcal{S}$ using a PointNet architecture [42].
To train $E_{points}$, we take advantage of paired input data, for which we are given both the point cloud $P$ and its corresponding L-String $l$. Passing $l$ through the recursive L-String encoder produces a latent point $s$. By applying $E_{points}$ on $P$, we obtain $\hat{s}$. The training optimizes the loss function
$L_{points}=\sum_{j}\left(\hat{s}_{j}-s_{j}\right)^{2},$ (4)
where $j$ loops over all samples in the training set. This loss encourages the point cloud encoder function $E_{points}$ to represent $P$ at the location $\hat{s}\in\mathcal{S}$ that represents its corresponding L-String $s$.
4.4 Inferring L-Strings from Input Point Clouds
At inference, the input is an unstructured point cloud $P$. This input is encoded in $\mathcal{S}$ using the point cloud encoder, and the resulting point is decoded using the recursive decoder as $l=D(E_{points}(P))$.
Errors in the predicted module parameters of the reconstructed L-String can have a propagating effect on the global geometry of the plant reconstruction. For example, errors in predicting the angle of a stem that is located in the bottom of the plant can lead to deviation in the plant growth axis along the main stem. To avoid such deviations, we align the reconstructed plant with the input in an optimization framework. We optimize on certain parameters of the stem modules starting from the ones at the bottom of the plant and going up. This optimization is done on the angle and length parameters for the Chenopodium album plants, minimizing the one-sided Chamfer Distance from the reconstructed stem to the input point cloud. We then optimize on the parameters of the individual leaf modules that define the leaf size and curvature. The latter optimization is applied on the leaves in no particular order, minimizing the single directional Chamfer Distance between the reconstructed plant and the input point cloud. This results in a parametric L-String representation $l$ that allows for various downstream phenotyping tasks. In this paper, we focus on the following tasks.
3D Reconstruction
can be solved directly by applying the geometric interpretation rules on $l$ to retrieve the virtual plant 3D scene.
Skeletonization
is solved by applying geometric interpretation rules drawing organ skeletons on $l$.
Segmentation
is solved by applying geometric interpretation rules on $l$ and by keeping the labels of the organ types. The labels are propagated to $P$ by assigning each point in $P$ the most frequently assigned label among its $k$ nearest neighbors in the annotated point cloud corresponding to $l$.
5 Dataset
To train and evaluate our method, we design and generate a synthetic dataset of L-String and point cloud pairs of the Chenopodium Album plant. First, we generate the L-Strings using LPy platform [2], with L-system production and geometric interpretation rules, that we optimize to generate realistic Chenopodium virtual plants. We define different time functions for the different plant parameters that guide the plant growth for a range of $[12,19]$ days to get Chenopodium plants in early stages. Then, we use the labeled points sampling method from [4], to obtain the corresponding point clouds.
The L-strings of the generated Chenopodium Album plants consist of 5 different modules: stem, cotyledon, petiole, leaf and branch. For each module, there is a different set of parameters that describe the organ represented by the module. The modules’ parameters are as follows:
  • Stem: diameter, length, growing angle and phyllotaxis angle.
  • Cotyledon: angle, length, nerve curvature factor.
  • Petiole: starting diameter, ending diameter , angle, length, and elasticity factor.
  • Leaf: nerve curvature factor, length and width.
  • Branch: branching angle and elasticity factor.
The dataset contains plants of different shapes and structures, in a way that for each different structure there is a number of plants of different shapes. We ensure the balance between the different structures in the dataset by fixing the number of different plants for each structure, in order to avoid any bias towards certain structures. We generate plants of 33 different structures with 100 different plants for each structure, resulting in a dataset of 3300 pairs of L-Strings and point clouds.The dataset is split into training ($80\%$), validation($10\%$) and test($10\%$) sets. Figure 5 shows examples of generated plants from the dataset.
We evaluate our method on a test set of point clouds of synthetic plants, that have structures seen in the training set. Additionally, we evaluate on the same test set with simulated Gaussian noise, and on monocular depth images of the set, to show our models robustness to acquisition noise and incomplete inputs.
Finally we make first tests of our method on 5 scans of real Chenopodium Album plants from [33], to test its ability to generalize to real data.
Figure 5: Examples of Chenopodium Album plants from our dataset.
6 Evaluation
This section presents the evaluation of our method with respect to three common phenotyping applications, namely 3D reconstruction, 3D skeleton extraction, and segmentation. For each application, we outline an evaluation protocol and compare to strong baselines.
6.1 Implementation Details
We implemented our method using Pytorch on Quadro RTX 5000 GPU. We learn a latent space $\mathcal{S}$ of dimension $dim_{\mathcal{S}}=100$ with the recursive encoder-decoder pair architecture in Fig 2 of hidden layer size $h_{recur}=200$, and the classifiers with a hidden layer $h_{class}=200$. We optimize the weights of the network using Adam optimizer with a learning rate of 0.001, decaying by factor of 0.5 every 50 epochs. For the point cloud encoder we train a PointNet point regression network with Adam optimiser on a learning rate of 0.001. The weights of the layers of all the trained networks are initialized using Xavier uniform initialization.
Testset Method Accuracy Completeness Hausdorff Silhouette # Comp.
Synthetic SIREN $\mathbf{0.0035\pm 0.024}$ $\mathbf{0.001\pm 0.0004}$ $\mathbf{0.165\pm 0.204}$ $0.0395\pm 0.01$ $34.1\pm 19.2$
Ours $0.0148\pm 0.019$ $0.0186\pm 0.021$ $0.595\pm 0.215$ $\mathbf{0.0203\pm 0.008}$ $\mathbf{1}$
Noise SIREN $0.0156\pm 0.076$ $\mathbf{0.0019\pm 0.0006}$ $\mathbf{0.324\pm 0.327}$ $0.1425\pm 0.032$ $256.8\pm 69.4$
Ours $\mathbf{0.014\pm 0.017}$ $0.0182\pm 0.0186$ $0.598\pm 0.216$ $\mathbf{0.0204\pm 0.008}$ $\mathbf{1}$
Depth images SIREN $\mathbf{0.0098\pm 0.123}$ $\mathbf{0.0027\pm 0.0029}$ $\mathbf{0.428\pm 0.437}$ $0.0485\pm 0.03$ $41.2\pm 20.5$
Ours $0.0165\pm 0.02$ $0.021\pm 0.035$ $0.615\pm 0.22$ $\mathbf{0.0208\pm 0.008}$ $\mathbf{1}$
Table 2: Comparison to SIREN [44] for the 3D reconstruction of plants from point clouds on our test sets of synthetic scans, noisy point clouds and depth images. We compare on geometric distance measures (accuracy, completeness and Hausdorff), and on perceptual and qualitative measures (silhouette error and the number of connected components, denoted # Comp.).
Our method works on binary trees, while the L-String trees in our dataset mostly don’t have a binary tree structure. We summarize the L-Strings into binary trees, by combining the modules that always occur together as individual nodes with concatenated parameter sets. In particular, we combine the two cotyledon modules into one Cotyledons node with 4 parameters. Stems are always followed by a petiole and a leaf, therefore we combine them in one node called Stem that has 12 parameters. In case of a branching, the branch module is combined with the stem, petiole and leaf modules to form a Branch node of 14 parameters. Finally, the first stem module of the tree forms a node on its own called Root of 3 parameters. Note that we do not duplicate the symmetric parameters in the cotyledons (first two), which explains the number of resulting parameters. As for the Root node, we do not include the angle parameter since it is fixed across all the plants.
For the segmentation experiment where we use the $k$ nearest neighbors algorithm, we set $k=10$. As for the reconstruction experiments, when using SIREN we need to input the point normals along with the point cloud, for this we use a normal estimation PCA-based method, which fits local planes to each point’s neighborhood.
6.2 3D Reconstruction
We quantitatively evaluate our method for 3D reconstruction of plants from input point clouds, on our test sets using five evaluation metrics. The first three are the commonly used metrics for 3D reconstruction methods, namely accuracy, completeness, and Hausdorff distance. Accuracy and completeness are computed using the Chamfer distance between the ground truth point cloud and uniformly sampled points on the reconstructed surface. Accuracy measures how closely the reconstructed points lie to the ground truth, while completeness measures how well the reconstruction covers all parts of the ground truth’s geometry. Hausdorff distance measures the largest distances between the reconstructed and ground truth points, offering a worst-case assessment of the geometric alignment.
As these metrics are only based on 3D distances and do not account for the visual consistency of the reconstructions, we additionally consider the number of connected components that are reconstructed and silhouette error. The silhouette error is computed as the average pixel-wise absolute difference between the binary masks of the reconstructed and ground-truth meshes, rendered from multiple viewpoints. For each viewpoint, both meshes are rendered to produce binary images (silhouettes), where foreground pixels indicate visible parts of the shape. The error is then measured by taking the absolute value of the pixel-wise subtraction between the two silhouettes, averaged across all pixels and viewpoints (we use 400 viewpoints). This metric evaluates the visual similarity between the ground truth and the reconstructed mesh. Finally, the number of connected components measures global consistency and should be $1$ for our dataset of isolated plants.
Figure 6: Qualitative results for 3D reconstruction by our method and SIREN [44], on synthetic clean, noisy and partial point clouds.
We compare our method to SIREN [44], an implicit neural representation method that uses sine activation functions to model continuous signals, including signed distance functions (SDFs). For 3D reconstruction, the surface can be extracted as the zero level set of the SDF using Marching Cubes. SIREN was identified as the best-performing method for reconstructing plant geometry from point clouds by Prasad et al. [38], due to its ability to capture fine geometric details.
Our method produces a complete smooth mesh that captures the plant visual shape and is robust with respect to input noise and missing input. Table 2 shows the results, where one can note that SIREN performs well in terms of the 3D distance-based error measures accuracy, completeness and Hausdorff distance. SIREN’s performance degrades significantly for depth images, and even more for noisy data. While our method has higher numerical error for these measurements than SIREN, it is much more robust to partial observations and noise. When considering silhouette loss, our method significantly outperforms SIREN as the visual similarity of the results is much higher for our method, (see Fig. 6). While the number of connected components of the reconstruction is always $1$ for our method by construction, SIREN reconstructs highly fragmented shapes.
Fig. 6 shows examples of SIREN’s and our reconstructions for the three types of test examples. Note that while our reconstructions may have some misalignments, they are well structured and similar to the ground truth. In contrast, SIREN reconstructs fragmented shapes close to the input point clouds that are globally dissimilar from the ground truth plants.
6.3 Skeletonization
Synthetic Noise Depth images
[53] $0.561\pm 0.312$ $0.701\pm 0.344$ $0.595\pm 0.324$
[6] $0.544\pm 0.286$ $0.583\pm 0.301$ $0.716\pm 0.377$
[30] $0.740\pm 0.355$ $0.776\pm 0.363$
Ours $\mathbf{0.037\pm 0.052}$ $\mathbf{0.034\pm 0.044}$ $\mathbf{0.042\pm 0.055}$
Table 3: Comparison to Xu et al. [53], Chaudhury et al. [6], and Livny et al. [30] for 3D skeleton extraction in terms of Chamfer Distance. “-” means that the method crashed due to numerical problems.
We quantitatively evaluate our method for 3D skeleton extraction from input point cloud using the bidirectional Chamfer Distance. Chamfer Distance helps capturing the geometrical discrepancy between the ground truth skeleton and the predicted skeleton by measuring the sum of nearest neighbour distances in both directions.
We compare our method with two classical plant skeletonization baselines Livny et al.[30], Xu et al. [53] and a stochastic skeleton refinement method Chaudhury et al. [6] that takes a predicted skeleton as input and outputs a refined skeleton. In this comparison we apply the skeleton refinement on the output skeletons from applying Xu et al. [53] on the plant point clouds.
Figure 7: Qualitative results for skeletonization by our method, Livny et al. [30], Xu et al. [53], and Chaudhury et al. [6] refinement on Xu skeleton, from clean synthetic, noisy and partial point clouds taken from monocular depth maps. “-” means that the method crashed due to numerical problems.
All these methods are designed to take scans of leafless plants as input. However, our dataset contains Chenopodium Album plants with leaves. Table 3 shows the results on the test sets of synthetic clean, noisy and partial point clouds. Our method shows consistent performance on the different test sets, suggesting robustness to noise and missing parts in the input point clouds for the skeletonization task. Since the baseline methods are not designed to take leaves as input, they tend to interpret leaf geometry as part of the skeleton, while the ground truth skeleton mainly describes the skeletal structure of the stem, branches and petioles. This explains the higher Chamfer distances, and the visual difference shown in Fig. 7.
6.4 Segmentation
Stem Leaf Petiole Cotyledons
Prec. Rec. F1 IoU Prec. Rec. F1 IoU Prec. Rec. F1 IoU Prec. Rec. F1 IoU
Synthetic
PlantNet 91 97 94 88 98 99 99 97 89 76 82 70 97 94 95 91
PSegNet 91 99.5 95 90 99 99.6 99 98 95 88 92 85 99 89 94 88
Ours 89 93 91 84 97 98 98 95 68 54 59 45 97 98 97 94
Noise
PlantNet 88 81 84 73 93 99 96 91 75 32 45 29 74 89 80 68
PSegNet 99 75 85 74 90 99.7 95 90 76 21 33 20 98 80 88 79
Ours 85 90 88 78 97 97 97 94 60 47 51 37 93 89 91 84
Table 4: Comparison to PlantNet [26] and PSegNet [25] for semantic segmentation on synthetic and noisy points test sets using standard classification measures Precision, Recall, F1-Score and Intersection over Union (IoU). All values are percentages.
We quantitatively evaluate our method’s performance on the semantic segmentation task applied to our test sets using standard classification metrics: precision, recall, F1 score, and Intersection over Union (IoU). For each class, precision is defined as the ratio of true positives to the sum of true positives and false positives, while recall is the ratio of true positives to the sum of true positives and false negatives. The F1 score is the harmonic mean of precision and recall. Finally, IoU is computed as the ratio of the intersection to the union of the predicted and ground truth point sets for a given class.
We compare our method to the two strong baselines PlantNet [26] and PSegNet [25]. Both PlantNet and PSegNet take a deep neural network approach with special sampling strategies, and are suitable for semantic and instance segmentation of plants.
Table 4 shows the results on the test sets of synthetic and noisy point clouds. We do not test segmentation methods on the depth image test set, as we do not have ground truth labels in this case. For synthetic scans, PSegNet outperforms all the methods except on Cotyledons segmentation, where our method shows a better performance. However, when noise is introduced, our method outperforms the other methods for most of the measures. Note that the petiole is the most difficult part to segment; this is because of its cylindrical shape, which is very close to the stem branches and makes it difficult to discriminate between these two classes. Overall, our method is robust to noise and performs on-par with the strong baselines.
Fig. 8 shows the semantic segmentation outputs of our method compared to the ground truth, where the semantics are properly captured even for noisy inputs. Furthermore, we show our method’s ability to perform instance segmentation in Fig. 9, where it successfully separates individual plant parts and remains robust to noise.
Figure 8: Examples of semantic segmentation results on synthetic clean and noisy point clouds.
Figure 9: Examples of instance segmentation results on synthetic clean and noisy point clouds.
7 Conclusion and Discussion
In this paper we have presented a data-driven method to infer plants parametric representations i.e. L-Strings, from 3D unstructured point cloud input. We have shown that the L-String representation contains structural and geometric information that the input scans lack, and that such information allows for multiple tasks like 3D reconstruction, skeletonization and segmentation. Our results have shown that our method is on-par with the state-of-the-art for these tasks on synthetic data, and is robust to noisy and incomplete input. However, the performance of our method is limited to synthetic data, and first results from this first version of the method need to be qualitatively and quantitatively improved when applied to real scans.
Figure 10: Our 3D reconstruction results on input scans of real Chenopodium Album plants.
This work demonstrate the feasibility to learn latent space of 3D plant structures and reconstruct plant architecture parametric models from it. In the next step, our aim is to upscale our method to address fully real 3D plant point clouds. For this we need to investigate the level of accuracy required to simulated plants real plants so that the method can use simulated plants to learn real plant architectures. Fig 10 shows first tests of our 3D reconstructions from real scans. Although the reconstructions are not yet accurate on all levels, global information is learned by our model.
8 Acknowledgements
We thank Franck Hétroy-Wheeler for helpful discussions, and Ayan Chaudhury, Frédéric Boudon, and Aymen Merrouche for providing support for their code. This work was partially supported by French government funding managed by the National Research Agency under grant ANR-24-CE23-1586 (4DPlants).

### v2

@@
 Abstract
-Reconstructing faithfully the 3D architecture of plants from unstructured observations is a challenging task.
-Plants frequently contain numerous organs, organized in branching systems in more or less complex spatial networks, leading to specific computational issues due to self-occlusion or spatial proximity between organs.
-Existing works either consider inverse modeling where the aim is to recover the procedural rules that allow to simulate virtual plants, or focus on specific tasks such as segmentation or skeletonization.
-We propose a unified approach that, given a 3D scan of a plant, allows to infer a parameterized representation of the plant.
-This representation describes the plant’s branching structure, contains parametric information for each plant organ, and can therefore be used directly in a variety of tasks.
-In this data-driven approach, we train a recursive neural network with virtual plants generated using an L-systems-based procedural model.
+Plants frequently contain numerous organs, organized in 3D branching systems defining the plant’s architecture.
+Reconstructing the architecture of plants from unstructured observations is challenging because of self-occlusion and spatial proximity between organs, which are often thin structures.
+To achieve the challenging task, we propose an approach that allows to infer a parameterized representation of the plant’s architecture from a given 3D scan of a plant.
+In addition to the plant’s branching structure, this representation contains parametric information for each plant organ, and can therefore be used directly in a variety of tasks.
+In this data-driven approach, we train a recursive neural network with virtual plants generated using a procedural model.
 After training, the network allows to infer a parametric tree-like representation based on an input 3D point cloud.
 Our method is applicable to any plant that can be represented as binary axial tree.
-We evaluate our approach on Chenopodium Album plants, using experiments on synthetic plants to show that our unified framework allows for different tasks including reconstruction, segmentation and skeletonization, while achieving results on-par with state-of-the-art for each task.
+We quantitatively evaluate our approach on Chenopodium Album plants on reconstruction, segmentation and skeletonization, which are important problems in plant phenotyping.
+In addition to carrying out several tasks at once, our method achieves results on-par with strong baselines for each task.
+We apply our method, trained exclusively on synthetic data, to 3D scans and show that it generalizes well.
 Figure 1: Our method takes a 3D point cloud of a plant as input, and outputs a parameterized representation of the plant.
 This representation encodes the plant’s branching structure and geometry along with semantic information such as organ type, and allows for multiple tasks including reconstruction, organ segmentation and skeleton extraction.
-11footnotetext: Inria centre at the University Grenoble Alpes22footnotetext: Inria centre of Lyon
 1 Introduction
-Inferring high-level information of plants from observations in a way that respects the biological traits (e.g.
-the radius of a lateral branch is never larger than that of it’s parent stem) is an important problem that has been studied for decades [52].
-High-level information includes the reconstruction of the 3D geometry of a plant, the segmentation of a plant into parts, and the extraction of a skeleton or measurements.
+Plant phenotyping consists of quantifying how plant’s genotypes grow in their environment and has important applications in crop management.
+To automate this task, it is necessary to infer high-level information of plants from observations.
+This problem, which has been studied for decades [phenotyping2015], can be decomposed into a number of subtasks.
+High-level information includes the reconstruction of the 3D geometry of a plant, the segmentation of a plant into parts, and the extraction of a skeleton.
 Observations are typically 2D images or 3D scans.
-Automatically extracting this information allows to use reconstructed 3D plants in content generation and virtual reality [29], and to use part segmentation, skeleton information, and measurements in plant phenotyping, e.g.
-[33, 13, 18].
-The problem of inferring high-level information is challenging because plants have complex structures due to their branching systems, which leads to strong self-occlusions and ambiguities.
+Automatically extracting this information has applications in plant phenotyping, e.g.
+[mirande2022graph, dobbs2023smart, ghahremani2021direct], and to use reconstructed 3D plants in content generation and virtual reality [TreePartNet21].
+Such an inference task is challenging as plants have complex structures due to their branching systems, which lead to strong self-occlusions and ambiguities.
 Existing works that analyze 2D or 3D plant observations therefore either focus on inverse modeling, where the goal is to find growth rules that allow to generate a given plant, or are task-specific.
-Inverse modeling is challenging and existing methods are limited to simple branching structures [24, 48].
+Inverse modeling is challenging and existing methods are limited to leafless branching structures [guo2020inverse, stava2014inverse].
 Task-specific methods focus on 3D reconstruction e.g.
-[29], organ segmentation e.g.
-[33], skeletonization e.g.
-[13], or measurement extraction e.g.
-[18].
-We propose a unified approach that, given a 3D scan of a plant, allows to infer a parameterized representation of the plant.
-This representation contains parametric information on each plant organ, and can therefore be directly used to reconstruct the 3D geometry of the plant with corresponding segmentation labels and measurements, thereby allowing to solve a variety of tasks.
-We achieve this by learning a shape space of 3D plants based on representations generated by a biologically inspired procedural model, in a way that for any point in shape space, we can retrieve its corresponding parameterized representation.
-By learning a function that maps an observed 3D point cloud to the learned shape space, our work allows to find the representation of a plant without reconstructing the full procedural model rules.
-Hence, our work can be considered a step towards inverse modeling-based reconstruction.
-The key idea that allows our method to address the challenging problem of inferring high-level information from 3D scan data plagued by occlusions is to learn a shape space with the help of instances generated using a procedural model made of recursive rules that describe the development of the plant architecture.
-We leverage a Lindenmayer-systems (L-systems) model [40] allowing to generate plants, and the recursive nature of plant growth in a recursive neural network to learn a latent distribution of the plants’ geometry and branching structure.
-We generate a synthetic dataset using procedural modeling rules, which allows for supervised learning from large amounts of occlusion-free training data.
-The synthetic plants are represented in a binary axial tree form that can be geometrically interpreted to obtain the 3D virtual plant.
-The binary tree format is exploited using a recursive auto-encoder.
-To address the problem of generalizing to 3D scans plagued by acquisition noise and occlusions, we virtually simulate the acquisition noise on the 3D virtual plants, in addition to simulated monocular depth maps during training and learn a mapping from simulated 3D scan data to the shape space.
-We evaluate our method on synthetic Chenopodium Album plants, an annual plant, in early growth stages.
-We choose this plant because of the availability of real scans in dataset [34], and its simple structure in early stage that can be encoded into a binary tree representation.
-We use our method to perform the different tasks of 3D reconstruction, skeleton extraction and plant part segmentation.
-Our experiments show results that are on-par with the state-of-the-art, with more robustness to noisy and partial inputs.
-We also show qualitative results for instance segmentation, and discuss the potentials to apply our method on real scans.
+[TreePartNet21], organ segmentation e.g.
+[mirande2022graph], or skeletonization e.g.
+[dobbs2023smart].
+We propose a method to infer a parameterized representation given as input a 3D scan of a plant.
+This is achieved by learning a shape space of 3D plants that captures both the plant’s structure and the parametric 3D shape of all plant organs.
+The resulting representation can be directly used to reconstruct the 3D geometry of the plant with labels.
+This offers the key advantage of allowing to solve a variety of tasks, while allowing for complex plant architectures.
+Learning shape spaces of plants is under-explored due to challenges arising from variation in both plant structure and organ shape.
+A recent work proposes a first solution [cheng2025demeter], while relying on large captured annotated datasets for training and requiring pre-segmented input for inference, both of which are costly and error-prone.
+In contrast, we propose a method trained purely on synthetic data and show its applicability to raw 3D scan acquisitions without annotations.
+To learn a shape space of 3D plants, we leverage biologically inspired procedural models made of recursive rules that describe plant development.
+Procedural models inform both the design of our neural network that learns a shape space of 3D plants and the generation of large amounts of synthetic training data.
+We use Lindenmayer-systems (L-systems) to represent plants in a binary axial tree form [Prusinkiewicz1990TheAB].
+In L-systems, each plant architecture is represented as a bracketed string of parameterized modules, the L-String.
+We use a recursive neural network to model the recursive nature of the L-string.
+Without loss of generality, we restrict plant representations to binary trees, which allows the use of recursive auto-encoders trained on (binary) L-Strings.
+In this way, we learn both branching structure and shape distribution of plant organs from synthetic data.
+To generalize to acquired 3D scans of plants, we learn a mapping from simulated 3D scan data to the learned shape space.
+For robustness, we simulate acquisition noise on the virtual plants.
+We evaluate our method on Chenopodium album plants, an annual plant, in early growth stages.
+Chenopodium is an ideal plant for our tests as it is a small plant (as opposed to trees) while still displaying a complex branching architecture and has small leaves that partially occlude each other, but not excessively.
+In addition, a collection of real scans are readily available at different stages of development [mirande_2022_6962994].
+We thus created for this plant a large dataset of 3D virtual plants represented as L-Strings, with associated 3D point clouds simulated with different types of acquisition noise.
+We use our method to perform 3D reconstruction, skeleton extraction and plant organ segmentation. Our experiments show results on-par with strong baselines, and robust to noisy and partial inputs.
 The main contributions of this work are:
-  • A unified approach to infer a parameterized representation of a plant from a 3D scan based on a recursive neural network.
-  • 3D plant reconstruction, skeletonization, organ segmentation, with performance that is on-par with the state-of-the-art on each task.
-  • A dataset of virtual plants, in a form of L-Strings, that represents instances of Chenopodium Album plants in early growth stages.
+  • An approach to infer a parameterized representation of a plant from a 3D scan based on a shape space of 3D plants learned using a recursive neural network.
+  • 3D plant reconstruction, skeletonization and segmentation, with performance on-par with strong baselines.
+  • A dataset of virtual plants, in the form of L-Strings, that represents instances of Chenopodium Album plants in early growth stages.
+  • A demonstration that learning based exclusively on virtual plants can be generalized to fit real plants.
 2 Related Work
-Vision based systems and platforms have been developed for plant phenotyping and crops analysis [7, 8, 11, 10, 37].
-Furthermore, plant scan datasets and data generation methods were provided for training data-driven models, e.g.
-the rosebush annotated dataset ROSE-X [15], an annotated dataset of 3 plant species PLANEST-3D [31], and a labeled point cloud generation method of virtual plants [4].
-Furthermore, small datasets of 3D scans were captured e.g.
-[34].
-Existing works on plant phenotyping that deal with plant scan data belong to two main categories: works that aim for inverse plant modeling, and task-specific methods that aim to extract information.
-Our work is in between these categories, since it is centered around the inference of parameterized plant representations, which allows to extract various high-level information including the full branching structure and 3D geometry of the plant.
-This can be seen as a step towards inverse plant modeling as we infer an instance of a procedural model.
+Existing works to infer high-level information of plants belong to two main categories: inverse plant modeling and task-specific methods.
+Our work is between these categories, since it infers a parameterized plant representation, which includes the branching structure and 3D geometry of the plant.
+This infers a step of a procedural model and can be seen itself as a step towards inverse plant modeling.
+A notable exception to our categorization is Demeter [cheng2025demeter], a work close in spirit to ours.
+Demeter learns a parametric plant model and demonstrates its use for reconstruction tasks.
+Unlike our method, Demeter requires manually annotated 3D scans for training and uses an elaborate three-step method for reconstruction, which requires pre-segmented input.
+In contrast, our method is trained on virtual data, which is cheaper to obtain, and directly infers a parametric representation without costly data pre-processing.
 2.1 Inverse Plant Modeling
-Procedural plant models generate a plant following growth rules.
-On the contrary, inverse plant modeling aims to find the rules that allow to generate the given plant.
-In other words, the goal is to infer the model given the plant geometry as input.
-Guo et al.
-[23] and Štáva et al.
-[47] infer procedural modeling rules generated using an L-system from 2D images of simple branching structures.
-Štáva et al.
-[48] introduce a parametric procedural model and estimate the parameters that allow this model to generate trees of a single species similar to an input 3D mesh.
-More recently, Lee et al.
-[24] used transformers to learn a latent space that allows to predict L-system rules that correspond to 3D point clouds of a tree, and to generate similar trees.
-As Lee et al.
-[24], we use data-driven neural networks to learn a latent representation of the plant.
-However, we aim to directly predict the L-String that represents the input plant with its branching structure, geometrical and other biological traits, instead of predicting the L-system rules.
+Procedural models generate plants following growth rules.
+In contrast, inverse plant modeling aims to find the rules that allow to generate a given plant.
+Guo et al[guo2020inverse] and Štáva et al[vst2010inverse] infer procedural modeling rules generated using an L-system from 2D images of leafless branching structures.
+Štáva et al[stava2014inverse] introduce a parametric procedural model in 3D specific to trees.
+None of these methods is applicable to annual plants with leaves.
+More recently, Lee et al[10.1145/3627101] used transformers to learn a generative model to represent L-system rules, without the aim of reconstructing a precise instance of a given real plant input.
+CropCraft [zhai2024cropcraftinverseproceduralmodeling] performs inverse procedural modeling taking as input a collection of 2D images, and not 3D scans.
 2.2 Task Specific Methods
-Existing methods used in plant phenotyping from 3D input point clouds mainly focus on four tasks: 3D reconstruction of the plant, extracting a skeleton of the branching structure of the plant, segmenting the plant into organs, and extracting measurements from the point cloud.
+Work Recon- Skeletoni- Segment- Annual Direct
+struction zation ation plants infer-
+ence
+Gonzalez et al[gonzalez2023tree]
+Liu et al[TreePartNet21]
+Du et al[du2019adtree]
+Linvy et al[livny2010automatic]
+Preuksakarn et al[plant-recon]
+Parsad et al[prasad2022deep]
+Dobbs et al[dobbs2023smart]
+Chaudhury et al[Chaudhury2020.02.15.950519]
+Yan et al[5246837]
+Meyers et al[Meyer_2023]
+Mirande et al[mirande2022graph]
+Turgut et al[turgut2022segmentation]
+Turgut et al[TURGUT2022138]
+Wahabzada et al[2015article]
+Li et al[LI2022243]
+Li et al[psegnet2022]
+Cheng et al[cheng2025demeter]
+Ours
+Table 1: Positioning of our method w.r.t.
+the ability to perform 3D reconstruction, skeletonization and segmentation, applicability on annual plants, and ability to perform inference directly without requiring pre-segmentation of the input data.
+Existing plant phenotyping methods from 3D point clouds mainly focus on three tasks: 3D reconstruction, extracting a skeleton of the branching structure, and segmentation.
 3D Reconstruction
-Reconstructing plant geometry from an input scan represented by a 3D point cloud is an important problem due to its applications in computer graphics and plant phenotyping.
-This is a particularly challenging 3D reconstruction problem as plants contain thin structures and as different organs lead to significant self-occlusions.
-In computer graphics, Gonzalez et al.
-[22] use UnrealEngine [41] to reconstruct digital urban trees by reconstructing a mesh that represents the tree trunk, estimating the volume and density of the canopy, and filling the canopy with generated leaves.
-This method focuses on global geometry and ignores the branching structure of the plant architecture.
-Another line of work focuses on the tree branching structure [14, 39, 54, 29, 53], by estimating a skeletal structure of the tree.
-Given the skeleton, a mesh surface can be reconstructed over each branch.
-However, the foliage part of the trees is either excluded from the input, or reconstructed by randomly generating leaves to fill the canopy.
-Closest to our work, Prasad et al.
-[38] explore different implicit surface reconstruction algorithms, including classical and neural based ones, on 3D point clouds of a single plant.
-They demonstrate that this task is challenging due to fine geometric structures.
-Note that our method’s output contains more information than a 3D reconstruction, as our parametric output allows to infer information on the plant’s skeleton and organs.
-For 3D reconstruction, we compare our method experimentally to this work.
+While some works reconstruct 3D plants from a single 2D image [liu2025boxplant], we focus on reconstructing plant geometry from an input 3D point cloud.
+This 3D reconstruction problem is challenging as plants contain thin structures and self-occlusions.
+In computer graphics, Gonzalez et al[gonzalez2023tree] reconstruct urban trees by computing a mesh representing the tree trunk, estimating the volume and density of the canopy, and filling the canopy with generated leaves.
+The branching structure of the plant is ignored.
+Another line of work focuses on branching structures of trees [du2019adtree, plant-recon, 5246837, TreePartNet21, 10.1145/1289603.1289610] by first extracting a skeleton and subsequently reconstructing a mesh per branch.
+Foliage is not reconstructed.
+Closest to our work, Prasad et al[prasad2022deep] compare different implicit surface reconstruction algorithms on 3D point clouds of a single plant.
+We compare our method experimentally to the best performing method tested in this work.
 Skeletonization
-Extracting the skeletal structure of plants is a well-studied problem. One line of works, discussed above, extracts skeletons as one part of a pipeline to reconstruct the plant [14, 39, 54, 29, 53].
-Another line of works corresponds to optimization-based methods.
-Meyer et al.
-[32] use a point contraction algorithm to extract skeletons of leafless cherry trees.
-Works using graph-based approaches have shown to extract skeletal structures even in the presence of noise for trees [53, 30].
-A recent deep learning-based approach uses a sparse convolutional neural network to estimate the medial axis of a tree [13].
-Given a point cloud of a plant with an initial extracted skeleton, Chaudhury et al.
-[6] propose a stochastic optimization to refine the initial skeleton by estimating point transformations.
-All discussed methods require a leafless input point cloud to estimate the skeleton.
-When leaves are present in the input point cloud, these methods estimate skeletal structures inside of the leaves, thereby adding noise to the output.
-Since our method outputs a segmented parameterized representation of the plant, extracting the skeleton of a leafy input plant does not pose additional challenges.
-We experimentally compare our method to one successful approach from each category for which code is available, namely [53, 30, 6].
+One line of work, discussed above, extracts skeletons as one part of a pipeline to reconstruct the plant [du2019adtree, plant-recon, 5246837, TreePartNet21, 10.1145/1289603.1289610].
+Other approaches use optimization-based methods.
+Meyer et al[Meyer_2023] use a point contraction algorithm to extract skeletons of leafless trees.
+Graph-based approaches have shown to extract skeletal structures even in the presence of noise [10.1145/1289603.1289610, livny2010automatic].
+A recent deep learning-based approach estimates the medial axis of a tree [dobbs2023smart].
+Given a point cloud of a plant with an initial extracted skeleton, Chaudhury et al[Chaudhury2020.02.15.950519] refine the skeleton.
+All discussed methods require leafless input.
+When leaves are present, skeletal structures are estimated inside the leaves, resulting in noisy output.
+Our method allows to extract skeletons with or without leaves.
+We experimentally compare our method to one successful approach from each category for which code is available [10.1145/1289603.1289610, livny2010automatic, Chaudhury2020.02.15.950519].
 Segmentation
-Segmenting different organs from point clouds of plants is critical for phenotyping on both semantic and instance levels.
-Both learning and optimization-based techniques were proposed in the literature to segment a plant into organs.
-Mirande et al.
-[33] propose a graph-based optimization approach with botanical knowledge refinement for semantic and instance segmentation.
-Wahabzada et al.
-[51] present an unsupervised data-driven method using histogram clustering in a Euclidean space.
-For supervised learning methods, a benchmark for plant organ segmentation [49] has been released based on the ROSE-X dataset [15].
-More complex deep learning architectures designed for organ segmentation lead to accurate results [26, 50, 25].
-We compare our method experimentally to two of these successful approaches [26, 25].
-Measurement Extraction
-Extracting phenotypic measures like leaf area index or main stem height provides insight about plant and crop behavior in different conditions.
-Paproki et al.
-[36] propose a mesh processing technique for 3D plants phenotypic traits analysis from multi-view images.
-Chattopadhyay et al.
-[3] model and reconstruct apple trees with diameter estimation for the application of automatic dormant pruning.
-Ghahremani et al.
-[18] apply the RANSAC algorithm to model plant organs that fit an input point cloud to estimate phenotypic features like angles and diameters of leaves and branches.
+Both learning and optimization-based techniques can segment a plant into organs.
+Mirande et al[mirande2022graph] propose a graph-based optimization approach with botanical knowledge refinement for semantic and instance segmentation.
+Wahabzada et al[2015article] present an unsupervised data-driven method.
+For supervised learning methods, a benchmark for plant organ segmentation [turgut2022segmentation] has been released based on the ROSE-X dataset [Dutagaci2020ROSEXAA].
+More complex deep learning architectures designed for organ segmentation lead to accurate results [LI2022243, TURGUT2022138, psegnet2022].
+We compare our method to the two generally applicable approaches [LI2022243, psegnet2022].
 Positioning
-Our experiments focus on the three tasks of 3D reconstruction, skeletonization and segmentation.
-As shown in Table 1, only few works perform the three tasks of 3D reconstruction, skeletonization and segmentation, like [29, 54].
-However, these methods are designed to be applied on big trees with trunks and foliage, and are not applicable on small annual plants where the leaf size to stem diameter proportion is different.
-On annual plants, our method remains generic due to the learned parametric model that contains the information needed for different tasks.
-Work Recon- Skeletoni- Segment- Annual
-struction zation ation plants
-Gonzalez
- et al. [22]
-Liu
- et al. [29]
-Du
- et al. [14]
-Linvy
- et al. [30]
-Preuksakarn
- et al. [39]
-Parsad
- et al. [38]
-Dobbs
- et al. [13]
-Chaudhury
- et al. [6]
-Yan
- et al. [54]
-Meyers
- et al. [32]
-Mirande
- et al. [33]
-Turgut
- et al. [49]
-Turgut
- et al. [50]
-Wahabzada
- et al. [51]
-Li
- et al. [26]
-Li
- et al. [25]
-Paproki
- et al. [36]
-Ghahremani
- et al. [18]
-Ours
-Table 1: Positioning of our method w.r.t. the ability to perform 3D reconstruction, skeletonization and part segmentation, and the applicability on small annual plants.
+Table 1 shows that only few works perform 3D reconstruction, skeletonization and segmentation [TreePartNet21, 5246837, cheng2025demeter].
+Two of these [TreePartNet21, 5246837] are limited to big trees with trunks and foliage, and not applicable on small annual plants.
+The third method [cheng2025demeter] takes 2D images as input and performs segmentation as pre-processing, and is therefore not comparable to our method.
 3 Background
-We propose a method that learns a shape space of a plant species with the help of recursive neural networks.
-To learn this shape space, we rely on a large dataset of 3D plants with corresponding parametric representations that has been simulated with the help of plant models.
-This section provides background on plant modeling and on recursive neural networks.
+This section provides background on plant modeling and recursive neural networks.
 3.1 Plant Modeling
 Plant modeling aims to find a mathematical model that describes the complex geometry and growth rules for a species of plants.
-Moreover, plant models respect biological rules to allow realistic simulation of the plant’s appearance and behavior in different environmental conditions [5, 19].
-The importance of plant modeling comes from its applications in biological and agronomical studies.
-The synthetic models can be used to perform growth simulation experiments, or estimate biological traits and growth patterns of plants.
-This allows to study plants with fast and non-invasive methods.
-Procedural methods are among the most commonly used plant modeling approaches, including L-systems [40], the space colonization algorithm [43], and commercial tools such as SpeedTree.
-In our work, we use procedural plant models to generate data for training and evaluating our neural network.
-Our implementation uses an L-system-based procedural approach [40] to generate plant data, a classical rule based plant modeling technique, that is particularly convenient due to the availability of tools such as L-Py [2].
+Moreover, plant models respect biological rules to allow realistic simulation of the plant’s appearance and behavior in different environmental conditions [Godin.2005].
+Procedural methods are among the most commonly used plant modeling approaches, including L-systems [Prusinkiewicz1990TheAB], the space colonization algorithm [space], and commercial tools such as SpeedTree [speedtree].
+In our work, we use procedural plant models to generate training and test data.
+Our implementation uses an L-system-based procedural approach [Prusinkiewicz1990TheAB] to generate plant data, a classical rule based plant modeling technique, that is particularly convenient due to the availability of tools such as L-Py [boudon2012py].
 In L-systems, a plant is represented as a string of symbols, called L-String, possibly bearing geometrical or biological parameters.
-Similarly to natural or computer language grammars, an L-system consists of a set of rewriting rules that define how plant components represented by L-String symbols change as time proceeds, by specifying how symbols get replaced by combinations of other symbols [40, 21].
+Similarly to natural or computer language grammars, an L-system consists of a set of rewriting rules that define how plant components represented by L-String symbols change as time proceeds, by specifying how symbols get replaced by combinations of other symbols [Prusinkiewicz1990TheAB, Godin.2005].
 At each step, the rules replace the symbols in parallel, resulting in a new L-String representing the next plant state.
@@
 Opening and closing brackets indicate the start and the end of every branch in the plant.
-Two successive modules in an L-sting have a parent-child relationship.
+Two successive modules in an L-string have a parent-child relationship.
 Open brackets allow for a module to have more than one child, and for different children to have siblings relationships.
 However, each parent has at most one special child that corresponds to its successor on the same plant axis.
-Such L-Strings encode axial trees [40] that represent the plant’s architecture.
-Modules in L-Strings can have parameters that give information about a plant organ.
-This includes the organ’s age, order in the branch or geometric attributes (e.g.
-length, radius, angle, etc.).
-By applying geometric interpretation rules on an L-String using turtle geometry [1], a 3D scene of the plant can be generated.
-The virtual plant comprises of connected 3D shapes corresponding to the modules.
-In other words, an L-String can encode a plant’s geometry in addition to its architecture.
-Our L-system models are implemented using the L-Py [2] language, that provides an extension of the Python programming language to design procedural models using the L-systems formalism.
-L-Py allows to create and manipulate L-system rules, handle the information present in an L-String, and generate 3D virtual plants from L-Strings.
+Such L-Strings encode axial trees [Prusinkiewicz1990TheAB] that represent the plant’s architecture.
+Modules in L-Strings can have parameters that give information about a plant organ, which allows an L-String to encode a plant’s geometry in addition to its topology, defined as its structure.
 3.2 Recursive Neural Networks
-Recursive Neural Networks (RvNN) [12] are a class of deep neural networks that apply the same network architecture recursively on input that is structured as a tree graph to give a single representation in a latent space.
-The term recursive refers to the network architecture being applied to the output of step $i-1$ during step $i$.
-The recursivity of RvNNs allows to handle input of varying size, as the network is applied in a bottom-up fashion from the leaves to the root.
-These networks learn hierarchical dependencies and correlations between input components.
-Socher et al.
-[46, 45] introduced an RvNN that learns information on different domains, including natural language processing and natural scene image processing, by recursively parsing a tree graph with binary structure.
-Models based on RvNN have been applied to natural language processing [9], 3D indoor scene generation [28, 17], blood vessel synthesis [16], and 3D shape structure recovery [35], segmentation [55] and generation [27].
-Of particular interest for our work are recursive auto-encoders for binary trees, such as the ones used to generate 3D shapes of man-made objects [27].
+Recursive Neural Networks (RvNN) [costa2003towards] are deep neural networks that apply the same network architecture recursively on structured input.
+The term recursive refers to the network being applied to the output of step i1i-1 during step $i$.
+The recursivity of RvNNs allows to handle input of varying size, as the network is applied bottom-up from the leaves to the root.
+RvNNs have been applied to different domains [socher2010learning, socher2011parsing] including natural language processing [chuan-an-etal-2018-unified], 3D generation [li2019grains, 10018465, 10.1145/3072959.3073637], blood vessel synthesis [10.1007/978-3-031-43907-0_7], 3D shape structure recovery [Niu_2018_CVPR], and segmentation [Yu_2019_CVPR].
+Of particular interest for our work are recursive auto-encoders for binary trees [10.1145/3072959.3073637].
 A recursive auto-encoder learns on a binary tree structure, where all nodes can be represented in a latent space of dimension $dim_{\mathcal{S}}$.
-The encoder follows the tree structure and recursively merges pairs of inputs to form a new point in the same latent space until the full tree is represented as a single point in latent space.
+The encoder follows the tree structure and recursively merges pairs of inputs to form a new point in the same latent space until the full tree is represented as a single latent point.
 Inversely, the decoder recursively decodes a single point into two points in the same latent space until the full tree structure is decoded.
-Figure 2 shows an example of the network architecture of a recursive encoder-decoder pair.
-The encoder-decoder pair allows to learn correlations between continuous properties encoded in features along the tree.
-An RvNN typically contains different types of encoder-decoder pairs.
-This allows to implement different operations, such as merging or splitting different parts.
-During decoding, given a point in latent space, an algorithm is needed to decide which decoder to apply and when to stop the recursion.
-To achieve this, the encoders and decoders are trained with ground truth structural information.
-During encoding, the information of which encoder to apply is available.
-For decoding, the RvNN jointly learns a set of classifiers that is applied before each decoding step.
+4 Method
+Our goal is to infer an L-String $l$ from an input 3D scan of a plant represented as point cloud $P$.
+Learning a direct regression from the space of 3D point clouds to the space $\mathcal{L}$ of L-Strings is difficult, as point clouds have varying numbers of points and are unstructured, and as L-Strings vary both in discrete (i.e.
+number and type of modules) and continuous (i.e.
+values of the angle and length parameters) ways.
+To address this problem, we combine an RvNN learned on $\mathcal{L}$ with an encoder that maps a point cloud to a latent space $\mathcal{S}$, as shown in Fig.
+2.
+During training, we first learn the latent space $\mathcal{S}$ that allows to represent L-Strings of a fixed plant species, see Sec.
+4.1.
+Second, we train a neural network called point cloud encoder that maps an input point cloud $P$ to a point $s\in\mathcal{S}$, see Sec.
+4.2.
+During inference, the input point cloud $P$ is encoded in $s\in\mathcal{S}$ using the point cloud encoder, and subsequently decoded to an L-String using the recursive decoder, see Sec.
+4.3.
+Figure 2: Overview: our method learns a latent space $\mathcal{S}$, that allows the mapping of 3D point clouds to L-Strings.
+At inference, the point cloud is mapped to $\mathcal{S}$ using the point cloud encoder on the left, and the resulting latent point allows to reconstruct the corresponding L-String using the L-String decoder on the right.
+4.1 Representing L-Strings in Latent Space
+We aim to learn a latent space $\mathcal{S}$ that represents instances of a plant species.
+The branching structure of the L-Strings may not always correspond to a binary tree structure.
+We therefore first simplify the L-Strings by summarizing the information of modules that always occur together in the plant species.
+The definition of co-occurring modules is manually done once per species.
+The resulting combined modules are called nodes in the following, and we design the combination rules to guarantee a binary tree structure after combination.
+To simplify notations, we call the L-String with binary tree graph structure $l\in\mathcal{L}$ in the following.
+Our goal is to learn an encoder function $E:\mathcal{L}\to\mathcal{S}$, and a decoder function $D:\mathcal{S}\to\mathcal{L}$, such that $l\approx D(E(l)),\;\;\forall l\in\mathcal{L}$.
+Learning to encode different shape and structural information in latent points of fixed dimension is challenging.
+Inspired by [10.1145/3072959.3073637], we use an RvNN to learn the hierarchical relations between the modules of the L-String, leveraging the recursive nature of plant structures [Godin.20103zn].
+To achieve this, each node of the tree graph is represented individually as point in latent space, $s_{i}\in\mathcal{S}$, using a node auto-encoder.
+In a second step, each subtree of $l$ is represented in $\mathcal{S}$ using a recursive auto-encoder.
+This procedure is carried out recursively from the leaves to the root resulting in a latent point that represents the full $l$ in $\mathcal{S}$.
+4.1.1 Node Auto-encoders
+Each type of node has a different set of parameters, e.g.
+angles, widths, radii.
+This results in nodes that have different dimensionality in general, and that are not directly comparable.
+To allow nodes with different numbers of parameters to be used as input in an RvNN, we first map the information of each node to $\mathcal{S}$.
+To achieve this, a node encoder-decoder pair is learned for each node type.
+In the following, let $E_{node,i}$ and $D_{node,i}$ denote the encoder and decoder for each type of node, with $i=1,\ldots,N$ and $N$ the number of types of nodes.
+Both $E_{node,i}$ and $D_{node,i}$ consist of one fully connected linear layer, followed by a tanhtanh activation.
+4.1.2 Recursive Auto-encoders
+Figure 3: The two criteria to merge/split points in latent space shown on an example tree.
+Merging is applied recursively in a bottom-up manner until the whole tree is merged into one point, while the splitting performs the inverse operation.
+After applying node encoding, an L-String is represented as a binary tree where all nodes are individually represented as points in $\mathcal{S}$.
+It can then serve as input to an RvNN auto-encoder.
+To recursively merge or split nodes, we design encoder-decoder pairs based on the relationship of the input nodes in the tree.
+Two nodes to be merged can either be siblings or have a parent-child relationship in the tree.
+Based on this observation, we consider the two merging/splitting criteria shown in Fig.
+3.
+For each criterion, we learn one recursive encoder-decoder pair.
+Both the sibling encoder-decoder $(E_{sib},D_{sib})$ and the parent-child encoder-decoder $(E_{pc},D_{pc})$ are implemented as shown in Fig 4.
 Figure 4: The network architectures used for the recursive encoder and decoder.
@@
 $dim_{h}$ denotes the dimension of the hidden layer.
-L-Strings have tree-like hierarchical structure, and we use RvNNs to encode them into latent points while capturing the relationships between the different modules, and their features’ correlations along the tree.
-We also use RvNNs in the opposite direction, to retrieve the tree structure and its parameters.
-4 Method
-4.1 Overview
-We aim to obtain a tree-like parametric representation $l$ in the form of an L-String from an input point cloud $P$ that stems from an unstructured 3D scan of a plant.
-Learning a direct regression from the space of 3D point clouds to the space $\mathcal{L}$ of L-Strings is a difficult problem, as the point clouds have varying number of points and are unstructured, and as the L-Strings varies both in discrete (i.e.
-number and type of modules) and continuous (i.e.
-values of the parameters describing angles and lengths) ways.
-To address this problem, we combine an RvNN learned on $\mathcal{L}$ with an encoder that maps a point cloud to $\mathcal{S}$, as shown in Fig.
-3.
-During training, we first learn the latent space $\mathcal{S}$ that allows to represent L-Strings of a fixed plant species.
-This allows to represent 3D plants of varying shapes, with different number of organs, and different skeletal structures using latent representations $s\in\mathcal{S}$, and is explained in Sec.
-4.2.
-Second, we train a neural network called point cloud encoder that maps an input point cloud $P$ to a point $s\in\mathcal{S}$, as outlined in Sec.
-4.3).
-During inference, the input point cloud $P$ is encoded in the latent space as a point $s$ using the point cloud encoder, and subsequently decoded to an L-String using the recursive decoder, as detailed in Sec.
-4.4.
-Figure 3: Overview: our method learns a latent space $\mathcal{S}$, that allows the mapping of 3D point clouds to L-Strings.
-At inference, the point cloud is mapped to $\mathcal{S}$ using the point cloud encoder on the left, and the resulting latent point allows to reconstruct the corresponding L-String using the L-String decoder on the right.
-4.2 Representing L-Strings in Latent Space
-We aim to learn a latent space $\mathcal{S}$ that allows to represent various static instances of the same species of plant.
-Plants are represented in the space of L-Strings $\mathcal{L}$ and can be generated with a set of L-System rules.
-However, the branching structure of the L-Strings may not always correspond to a binary tree structure.
-We therefore first simplify the L-Strings by summarizing the information of modules that always occur together in the plant species.
-The resulting combined modules are called nodes in the following, and we design the combination rules to guarantee a binary tree structure after combination.
-To simplify notations, we call the simplified L-String $l\in\mathcal{L}$ in the following.
-Our goal is to learn an encoder function $E:\mathcal{L}\to\mathcal{S}$, and a decoder function $D:\mathcal{S}\to\mathcal{L}$, such that $l\approx D(E(l)),\;\;\forall l\in\mathcal{L}$.
-Learning to encode different shape and structural information in latent points of fixed dimension is challenging.
-Inspired by [27], we use an RvNN to learn the hierarchical relations between the modules of the L-String, leveraging the recursive nature of plant structures [20].
-To achieve this, each node of the tree is represented individually in a point in the latent space, $s_{i}\in\mathcal{S}$, using a node auto-encoder.
-In a second step, each subtree of the L-String is represented in $\mathcal{S}$ using a recursive auto-encoder.
-This procedure is carried out recursively from the tree-graph leaves up to its root.
-4.2.1 Node Auto-encoders
-Each type of node has a different set of parameters, e.g.
-angles, widths, radii.
-This results in nodes that have different dimensionality in general, and that are not directly comparable.
-To allow nodes with different number of parameters to be used as input in an RvNN, we first map the information of each node to $\mathcal{S}$.
-To achieve this, a node encoder-decoder pair is learned for each node type.
-In the following, let $E_{node,i}$ and $D_{node,i}$ denote the encoder and decoder for each type of node, with $i=1,\ldots,N$ and $N$ the number of types of nodes.
-Both $E_{node,i}$ and $D_{node,i}$ consist of a fully connected linear layer, followed by a $tanh$ activation.
-4.2.2 Recursive Auto-encoders
-Figure 4: The two criteria to merge/split points in latent space shown on an example tree.
-Merging is applied recursively in a bottom-up manner until the whole tree is merged into one point, while the splitting performs the inverse operation.
-After applying node encoding, an input L-String is represented as a binary tree where all nodes are individually represented as points in $\mathcal{S}$.
-It can then serve as input to an RvNN auto-encoder.
-To recursively merge or split the nodes, we design encoder-decoder pairs based on the relationship of the input nodes in the tree.
-Two nodes to be merged can be either siblings or have a parent-child relationship in the tree.
-Based on this observation, we consider the two merging/splitting criteria shown in Figure 4.
-For each criterion, we learn one recursive encoder-decoder pair.
-Both the sibling encoder-decoder $(E_{sib},D_{sib})$ and the parent-child encoder-decoder $(E_{pc},D_{pc})$ are implemented as shown in Fig 2.
 4.1.3 Auxiliary Classifiers
@@
 For the RvNN to decode structural information, it needs to learn to choose the appropriate decoder ($D_{sib}$, $D_{pc}$), or to end the recursive splitting.
-To predict the right decoder, we associate classes for the different splitting options (siblings, parent-child, stop) and jointly train a classifier on them.
-Due to different node types, the network should needs to learn which node decoder $E_{node,i}$ to apply to reconstruct each node of the L-string.
+To predict the right decoder, we associate classes for the different splitting options (siblings, parent-child, stop) and jointly train a classifier $C_{split}$ on them.
+The network needs to learn which node decoder $E_{node,i}$ to apply to reconstruct each node of the L-String.
 For that, a second classifier is trained, that allows to predict the node type for a latent point $s$ that is fully split.
-This classifier, called $C_{node}$ has $N$ output classes.
-Both classifiers share the architecture of a two-layer fully connected neural network.
-The first layer maps the input latent point to a hidden representation using a linear transformation followed by a $tanh$ activation.
-The second layer maps the hidden features to an output vector that corresponds to the different classes.
+This classifier is called $C_{node}$.
+Both classifiers use a two fully connected layers, with tanhtanh activation after the first layer.
 4.1.4 Training
@@
 For node $n$ in L-String $l$, with a set of parameters, the reconstruction loss $L_{rec}(n)$ is the mean squared error between the input node parameter vector that contains the plant part information and its reconstruction.
+The parameters’ weights are different in the loss.
 The reconstruction loss for the full L-String is
 $L_{rec}=\sum_{n}L_{rec}(n).$ (1)
-At each recursive encoding step $i$, the output point $s$ is a result of merging two points ($s_{1}^{(i)}$, $s_{2}^{(i)}$).
-Then in the inverse decoding step, $s$ will be split into ($\hat{s}_{1}^{(i)},\hat{s}_{2}^{(i)}$) that are supposed to represent the same information as ($s_{1}^{(i)}$, $s_{2}^{(i)}$) and should fall in the same region in $\mathcal{S}$.
-To learn a structured representation space into which we can map information from another modality (point clouds), we encourage corresponding latent points during encoding and decoding to be located close-by in $\mathcal{S}$ using a step-based recursion loss $L_{step}$ that is defined as
-$L_{step}=\sum_{i}\left(\left(s_{1}^{(i)}-\hat{s}_{1}^{(i)}\right)^{2}+\left(s_{2}^{(i)}-\hat{s}_{2}^{(i)}\right)^{2}\right).$ (2)
 The classifiers are trained with softmax classification and cross entropy loss.
@@
 Finally, the total loss that is used for training the recursive L-String auto-encoder is
-$L_{total}=L_{rec}+L_{step}+L_{split}+L_{node}.$ (3)
+$L_{total}=L_{rec}+L_{split}+L_{node}.$ (2)
 4.2 Point Cloud Encoder
-After representing L-Strings in latent space $\mathcal{S}$ has been learned, 3D plants can be represented as latent points $s\in\mathcal{S}$, and $s$ can be decoded into an L-String parametric representation $l$.
-To allow retrieving an L-String $l$ from an input point cloud $P$, we learn a new mapping function $E_{points}$ from the space of point clouds to $\mathcal{S}$ using a PointNet architecture [42].
-To train $E_{points}$, we take advantage of paired input data, for which we are given both the point cloud $P$ and its corresponding L-String $l$.
+After learning latent shape space $\mathcal{S}$, 3D plants can be represented as $s\in\mathcal{S}$, and $s$ can be decoded into a parametric L-String representation $l$.
+To infer an L-String $l$ from an input point cloud $P$, we learn a mapping function $E_{points}$ from the space of point clouds to $\mathcal{S}$ using a PointNet [DBLP:journals/corr/QiSMG16].
+To train $E_{points}$, we take advantage of paired input data, containing both the point cloud $P$ and its corresponding L-String $l$.
 Passing $l$ through the recursive L-String encoder produces a latent point $s$.
 By applying $E_{points}$ on $P$, we obtain $\hat{s}$.
-The training optimizes the loss function
+The training optimizes the loss
 $L_{points}=\sum_{j}\left(\hat{s}_{j}-s_{j}\right)^{2},$ (3)
-where $j$ loops over all samples in the training set.
-This loss encourages the point cloud encoder function $E_{points}$ to represent $P$ at the location $\hat{s}\in\mathcal{S}$ that represents its corresponding L-String $s$.
+where $j$ loops over all training samples.
+This loss encourages the point cloud encoder $E_{points}$ to represent $P$ at the location $\hat{s}\in\mathcal{S}$ that represents its corresponding L-String.
 4.3 Inferring L-Strings from Input Point Clouds
@@
 This input is encoded in $\mathcal{S}$ using the point cloud encoder, and the resulting point is decoded using the recursive decoder as $l=D(E_{points}(P))$.
-Errors in the predicted module parameters of the reconstructed L-String can have a propagating effect on the global geometry of the plant reconstruction.
+Errors in the predicted module parameters of the reconstructed L-String can lead to cumulated errors on the plant reconstruction.
 For example, errors in predicting the angle of a stem that is located in the bottom of the plant can lead to deviation in the plant growth axis along the main stem.
-To avoid such deviations, we align the reconstructed plant with the input in an optimization framework.
-We optimize on certain parameters of the stem modules starting from the ones at the bottom of the plant and going up.
-This optimization is done on the angle and length parameters for the Chenopodium album plants, minimizing the one-sided Chamfer Distance from the reconstructed stem to the input point cloud.
-We then optimize on the parameters of the individual leaf modules that define the leaf size and curvature.
-The latter optimization is applied on the leaves in no particular order, minimizing the single directional Chamfer Distance between the reconstructed plant and the input point cloud.
+To avoid such deviations, we align the reconstructed plant with the input in a test-time optimization framework.
+We optimize on parameters of the main stem modules starting from the ones at the bottom of the plant and going up.
+This optimization is done on 3D angle and length parameters.
+We then optimize on the parameters of the petiole modules that define the length and elasticity, and then on the parameters of leaf modules that define the leaf size and curvature.
+All the modules are optimized w.r.t.
+the bidirectional Chamfer Distance between the reconstructed plant and the input point cloud, and two optimization iterations from bottom to top are performed.
 This results in a parametric L-String representation $l$ that allows for various downstream phenotyping tasks.
 In this paper, we focus on the following tasks.
-3D Reconstruction
-can be solved directly by applying the geometric interpretation rules on $l$ to retrieve the virtual plant 3D scene.
-Skeletonization
-is solved by applying geometric interpretation rules drawing organ skeletons on $l$.
-Segmentation
-is solved by applying geometric interpretation rules on $l$ and by keeping the labels of the organ types.
+3D Reconstruction can be solved by applying the geometric interpretation rules on $l$ to retrieve the 3D plant.
+Skeletonization is solved by applying geometric interpretation rules on $l$ to reconstruct all stems and optionally main veins of leaves with minimal width.
+Segmentation is solved by applying geometric interpretation rules on $l$ and keeping the labels of the organ types.
 The labels are propagated to $P$ by assigning each point in $P$ the most frequently assigned label among its $k$ nearest neighbors in the annotated point cloud corresponding to $l$.
 5 Dataset
-To train and evaluate our method, we design and generate a synthetic dataset of L-String and point cloud pairs of the Chenopodium Album plant.
-First, we generate the L-Strings using LPy platform [2], with L-system production and geometric interpretation rules, that we optimize to generate realistic Chenopodium virtual plants.
-We define different time functions for the different plant parameters that guide the plant growth for a range of $[12,19]$ days to get Chenopodium plants in early stages.
-Then, we use the labeled points sampling method from [4], to obtain the corresponding point clouds.
+To train and evaluate our method, we design and generate a synthetic dataset of corresponding L-String and point cloud pairs of the Chenopodium Album plant.
+First, we generate the L-Strings using L-Py platform [boudon2012py], with L-system production and geometric interpretation rules, that we optimize to generate realistic Chenopodium virtual plants.
+We defined different time functions for the different plant parameters that guide the plant growth for a range of $[8,14]$ days to get Chenopodium plants in early stages.
+Then, we used the labeled points sampling method from [chaudhury20203d] to obtain the corresponding point clouds.
 The L-Strings of the generated Chenopodium Album plants consist of 5 different modules: stem, cotyledon, petiole, leaf and branch.
@@
 The modules’ parameters are as follows:
-  • Stem: diameter, length, growing angle and phyllotaxis angle.
+  • Stem: diameter, length, growing angle, bending angle and phyllotaxis angle.
   • Cotyledon: angle, length, nerve curvature factor.
@@
   • Branch: branching angle and elasticity factor.
-The dataset contains plants of different shapes and structures, in a way that for each different structure there is a number of plants of different shapes.
-We ensure the balance between the different structures in the dataset by fixing the number of different plants for each structure, in order to avoid any bias towards certain structures.
-We generate plants of 33 different structures with 100 different plants for each structure, resulting in a dataset of 3300 pairs of L-Strings and point clouds.The dataset is split into training ($80\%$), validation($10\%$) and test($10\%$) sets.
-Figure 5 shows examples of generated plants from the dataset.
-We evaluate our method on a test set of point clouds of synthetic plants, that have structures seen in the training set.
-Additionally, we evaluate on the same test set with simulated Gaussian noise, and on monocular depth images of the set, to show our models robustness to acquisition noise and incomplete inputs.
-Finally we make first tests of our method on 5 scans of real Chenopodium Album plants from [33], to test its ability to generalize to real data.
-Figure 5: Examples of Chenopodium Album plants from our dataset.
+The dataset contains plants of different shapes and structures.
+We balance the different structures in the dataset by fixing the number of different plants for each structure.
+We generated plants of 10 different structures with 100 different plants per structure, resulting in a dataset of 1000 pairs of L-Strings and point clouds.
+The dataset is split into training ($80\%$), validation ($10\%$) and test ($10\%$) sets.
+This dataset of realistic Chenopodium Album plants with ground truth parametric representation can serve as benchmark for plant reconstruction, segmentation and skeletonization tasks.
+We evaluate on the test set with clean point clouds, point clouds with simulated Gaussian noise, and monocular depth images of the set, to show our model’s robustness.
+To test generalization to real data, we test our method on 5 scans of real Chenopodium Album plants from Mirande et al[mirande_2022_6962994].
 6 Evaluation
-This section presents the evaluation of our method with respect to three common phenotyping applications, namely 3D reconstruction, 3D skeleton extraction, and segmentation.
+In this section we evaluate our method for the three common phenotyping applications 3D reconstruction, 3D skeleton extraction, and segmentation.
 For each application, we outline an evaluation protocol and compare to strong baselines.
-6.1 Implementation Details
+All methods are run on a Quadro RTX 5000 GPU.
+Implementation details are in appendix.
+Accuracy Completeness Size Time # Comp. LN Accuracy LAI Accuracy Topology
+Clean
+SIREN $\mathbf{0.0012}$ $\mathbf{0.0006}$ $780.88$ KB 7min 14s7\,min\,14\,s $33$
+Ours $0.0059$ $0.0090$ $\mathbf{17.56}$ KB $\mathbf{4\,min\,1\,s}$ $\mathbf{1}$ $\mathbf{98\%}$ $\mathbf{93\%}$ $\mathbf{75\%}$
+Noisy
+SIREN $0.0121$ $\mathbf{0.0008}$ $780.88$ KB 7min 27s7\,min\,27\,s $268$
+Ours $\mathbf{0.0054}$ $0.0074$ $\mathbf{17.56}$ KB $\mathbf{3\,min\,56\,s}$ $\mathbf{1}$ $\mathbf{98\%}$ $\mathbf{93\%}$ $\mathbf{78\%}$
+Depth maps
+SIREN $\mathbf{0.0013}$ $\mathbf{0.0022}$ $780.88$ KB 6min 44s6\,min\,44\,s $49$
+Ours $0.0060$ $0.0089$ $\mathbf{17.57}$ KB $\mathbf{3\,min\,47\,s}$ $\mathbf{1}$ $\mathbf{98\%}$ $\mathbf{94\%}$ $\mathbf{78\%}$
+Table 2: Comparison to SIREN [sitzmann2020implicitneuralrepresentationsperiodic] for 3D reconstruction on clean and noisy point clouds and depth maps.
+We compare geometric measures (accuracy, completeness), the model’s size and inference time, number of connected components of output (# Comp.), and phenotyping measures leaf number (LN) accuracy, leaf area index (LAI) accuracy, and topology.
+✗: the method cannot output the information.
+6.1 3D Reconstruction
+We quantitatively evaluate our method using eight complementary evaluation measures.
+The first two are accuracy and completeness, which are commonly used metrics assessing geometric alignment.
+Accuracy is the unidirectional Chamfer distance from the reconstructed surface to the ground truth, and completeness is the unidirectional Chamfer distance from the ground truth to the reconstructed surface.
+The next two measures are the output representation’s size and the inference’s running time, which measure the efficiency and compactness of the methods.
+We further report the number of connected components of the output meshes.
+We also assess the ability of models to predict correctly the number of leaves, defined as the mean percentage of matching leaf count between the reconstruction and the ground truth, and the mean percentage accuracy of leaf area index (LAI), defined as the one-sided leaf area per unit ground surface area [watson_lai].
+These two measures have been selected as they are commonly used in agronomy to calibrate crop models.
+Finally, topology accuracy is the percentage of reconstructed plants whose topology is correct (measured using a tree-edit distance [ferraro-hal-00827474]).
+Figure 5: Comparison to SIREN [sitzmann2020implicitneuralrepresentationsperiodic] for 3D reconstruction.
+We compare our method to SIREN [sitzmann2020implicitneuralrepresentationsperiodic], an implicit neural representation method that uses sine activation functions to model continuous signals, including signed distance functions (SDFs).
+For 3D reconstruction, the surface is extracted as the zero level set of the SDF using Marching Cubes.
+SIREN was identified as the best-performing method to reconstruct plant geometry from point clouds by Prasad et al[prasad2022deep].
+Table 2 shows the results.
+While SIREN performs well in terms of the 3D distance-based error measures accuracy and completeness for clean data, SIREN’s performance degrades significantly for noisy data.
+Our method is more robust w.r.t.
+noise and missing data, and outperforms SIREN in case of noise.
+Our representation is one to two orders of magnitude more compact than SIREN’s, and inference is twice as efficient.
+Unlike SIREN, which reconstructs a mesh without annotations, our method further allows to measure leaf number accuracy, LAI error and topology, and achieves very high accuracy on all these measures.
+Fig.
+5 shows 3D reconstructions for the three types of test examples.
+Note that our reconstructions are smooth, similar to the ground truth, and robust w.r.t.
+input noise and missing data.
+In contrast, due to its computational strategy, SIREN reconstructs fragmented shapes close to the input point clouds that are globally dissimilar from the ground truth, especially for noisy input.
+Finally, Fig.
+6 (second column) shows the 3D reconstructions obtained when applying our method to scans of real plants.
+Interestingly, despite local discrepancies, all plants are well reconstructed overall, which demonstrates that the methods generalizes well from virtual to real plant architecture data.
+Figure 6: Our method’s results on scans of real plants [mirande_2022_6962994]. Scans are reconstructed and segmented well overall, and the extracted skeletons closely follow the plants’ topologies.
+6.2 Skeletonization
+Full Branch
+Method Clean Noisy Depth images Clean
+[10.1145/1289603.1289610] $0.0102$ $0.0110$ $0.0145$
+[Chaudhury2020.02.15.950519] $0.0139$ $0.0154$ $0.0449$
+[livny2010automatic] $0.0257$ $0.0282$ -
+Ours $0.0178$ $0.0174$ $0.0199$ $0.0161$
+Table 3: Comparison for skeletonization w.r.t. Chamfer Distance. “-”: method crashed due to numerical problems. ✗: method’s scale cannot operate on the branch level.
+We quantitatively evaluate 3D skeleton extraction using the commonly used bidirectional Chamfer Distance.
+We compare our method to two classical plant skeletonization baselines Livny et al.
+[livny2010automatic], Xu et al[10.1145/1289603.1289610] and a stochastic skeleton refinement method Chaudhury et al[Chaudhury2020.02.15.950519] that takes a predicted skeleton as input and outputs a refined skeleton.
+The skeleton refinement is applied on the output skeletons of Xu et al[10.1145/1289603.1289610].
+All these methods are designed to take scans of leafless plants as input, and their output cannot be adjusted to different scales.
+In contrast, our method can output skeletons with or without the main veins of the leaves.
+Figure 7: Comparison to Xu et al[10.1145/1289603.1289610] for skeletonization on noisy input.
+Ours is applicable at different scales, and can output a full or a branch skeleton, while achieving visually accurate results.
+Table 3 shows the results for both the full skeleton that includes leaf veins and the branch skeleton that only includes the branching structure.
+Our method shows consistent performance on the different test sets, suggesting robustness to noise and missing parts in the input point clouds.
+Since baseline methods are not designed for input with leaves, they cannot output skeletons at different scales.
+Fig.
+7 shows a visual comparison on noisy input with the best method from table 3 Xu et al[10.1145/1289603.1289610], where our method achieves results close to the ground truth.
+More visual comparisons are in the appendix.
+Fig.
+6 (last column) shows skeletons extracted from scans of real plants.
+The skeletons closely follow the topology in the scans.
+6.3 Segmentation
+We compare our method to the strong baselines PlantNet [LI2022243] and PSegNet [psegnet2022].
+PSegNet provides labels on sparsely sampled point clouds, which we transfer to the full input point cloud using nearest neighbors similarly to what we do for our approach.
+Fig.
+8 shows the semantic segmentation results on examples from all test sets, where each type of plant organ is assigned a unique color.
+Our method is robust to noise and partial data, is on-par with the baselines, and outperforms PSegNet on the petioles segmentation.
+Quantitative comparisons on semantic segmentation and qualitative results on instance segmentation can be found in appendix.
+Fig.
+6 (third column) shows the segmentation of scans of real plants with our method.
+Note that all organs are well segmented overall.
+Figure 8: Comparison to PlantNet [LI2022243] and PSegNet [psegnet2022] for semantic segmentation on clean, noisy and partial point clouds.
+7 Conclusion
+In this paper we have presented a data-driven method to infer parametric representations of plants i.e.
+L-Strings, from 3D unstructured point cloud input.
+This advancement was possible by leveraging procedural models both for designing the neural networks to learn a parametric shape space for 3D plants, and for generating synthetic training data.
+We have shown that the L-String representation contains structural and geometric information that the input scans lack, and that such information allows for multiple tasks like 3D reconstruction, skeletonization and segmentation at once, with performance on-par with strong baselines for each individual task.
+Our results have further shown that our method, trained purely on synthetic data, generalizes well to noisy scans of real plants.
+8 Acknowledgements
+We thank Franck Hétroy-Wheeler for helpful discussions, and Ayan Chaudhury and Frédéric Boudon for sharing their codes.
+This work was partially supported by French government funding managed by the National Research Agency under grant ANR-24-CE23-1586 (4DPlants).
+Appendix
+9 Implementation Details
 We implemented our method using Pytorch on Quadro RTX 5000 GPU.
-We learn a latent space $\mathcal{S}$ of dimension $dim_{\mathcal{S}}=100$ with the recursive encoder-decoder pair architecture in Fig 2 of hidden layer size $h_{recur}=200$, and the classifiers with a hidden layer $h_{class}=200$.
+We learn a latent space $\mathcal{S}$ of dimension $dim_{\mathcal{S}}=64$ with the recursive encoder-decoder pair architecture in Fig.
+4 in the main paper, of hidden layer size $h_{recur}=128$, and the classifiers with a hidden layer $h_{class}=128$.
 We optimize the weights of the network using Adam optimizer with a learning rate of 0.001, decaying by factor of 0.5 every 50 epochs.
@@
 The weights of the layers of all the trained networks are initialized using Xavier uniform initialization.
-Testset Method Accuracy Completeness Hausdorff Silhouette # Comp.
-Synthetic SIREN $\mathbf{0.0035\pm 0.024}$ $\mathbf{0.001\pm 0.0004}$ $\mathbf{0.165\pm 0.204}$ $0.0395\pm 0.01$ $34.1\pm 19.2$
-Ours $0.0148\pm 0.019$ $0.0186\pm 0.021$ $0.595\pm 0.215$ $\mathbf{0.0203\pm 0.008}$ $\mathbf{1}$
-Noise SIREN $0.0156\pm 0.076$ $\mathbf{0.0019\pm 0.0006}$ $\mathbf{0.324\pm 0.327}$ $0.1425\pm 0.032$ $256.8\pm 69.4$
-Ours $\mathbf{0.014\pm 0.017}$ $0.0182\pm 0.0186$ $0.598\pm 0.216$ $\mathbf{0.0204\pm 0.008}$ $\mathbf{1}$
-Depth images SIREN $\mathbf{0.0098\pm 0.123}$ $\mathbf{0.0027\pm 0.0029}$ $\mathbf{0.428\pm 0.437}$ $0.0485\pm 0.03$ $41.2\pm 20.5$
-Ours $0.0165\pm 0.02$ $0.021\pm 0.035$ $0.615\pm 0.22$ $\mathbf{0.0208\pm 0.008}$ $\mathbf{1}$
-Table 2: Comparison to SIREN [44] for the 3D reconstruction of plants from point clouds on our test sets of synthetic scans, noisy point clouds and depth images.
-We compare on geometric distance measures (accuracy, completeness and Hausdorff), and on perceptual and qualitative measures (silhouette error and the number of connected components, denoted # Comp.).
-Our method works on binary trees, while the L-String trees in our dataset mostly don’t have a binary tree structure.
+Our method works on binary trees, while the L-String trees in our dataset do not have a binary tree structure in general.
 We summarize the L-Strings into binary trees, by combining the modules that always occur together as individual nodes with concatenated parameter sets.
-In particular, we combine the two cotyledon modules into one Cotyledons node with 4 parameters.
-Stems are always followed by a petiole and a leaf, therefore we combine them in one node called Stem that has 12 parameters.
-In case of a branching, the branch module is combined with the stem, petiole and leaf modules to form a Branch node of 14 parameters.
-Finally, the first stem module of the tree forms a node on its own called Root of 3 parameters.
-Note that we do not duplicate the symmetric parameters in the cotyledons (first two), which explains the number of resulting parameters.
-As for the Root node, we do not include the angle parameter since it is fixed across all the plants.
-For the segmentation experiment where we use the $k$ nearest neighbors algorithm, we set $k=10$.
+In particular, we combine the two cotyledon modules into one Cotyledons node with 6 parameters.
+Stems are always followed by a petiole and a leaf, therefore we combine them in one node called Stem that has 13 parameters.
+In case of a branching, the branch module is combined with the stem, petiole and leaf modules to form a Branch node of 15 parameters.
+Finally, the first stem module of the tree forms a node on its own called Root of 5 parameters.
+Parameters that are drawn from a Gaussian distribution do not correlate with their parent in the tree structure and cannot be learned by the recursive network, thus they are set to the mean during the training and are then optimized with other parameters in the test-time optimization phase.
+For the segmentation experiments where we use the $k$ nearest neighbors algorithm, we set $k=10$ for our results, and $k=3$ for PSegNet results.
 As for the reconstruction experiments, when using SIREN we need to input the point normals along with the point cloud, for this we use a normal estimation PCA-based method, which fits local planes to each point’s neighborhood.
-6.2 3D Reconstruction
-We quantitatively evaluate our method for 3D reconstruction of plants from input point clouds, on our test sets using five evaluation metrics.
-The first three are the commonly used metrics for 3D reconstruction methods, namely accuracy, completeness, and Hausdorff distance.
-Accuracy and completeness are computed using the Chamfer distance between the ground truth point cloud and uniformly sampled points on the reconstructed surface.
-Accuracy measures how closely the reconstructed points lie to the ground truth, while completeness measures how well the reconstruction covers all parts of the ground truth’s geometry.
-Hausdorff distance measures the largest distances between the reconstructed and ground truth points, offering a worst-case assessment of the geometric alignment.
-As these metrics are only based on 3D distances and do not account for the visual consistency of the reconstructions, we additionally consider the number of connected components that are reconstructed and silhouette error.
-The silhouette error is computed as the average pixel-wise absolute difference between the binary masks of the reconstructed and ground-truth meshes, rendered from multiple viewpoints.
-For each viewpoint, both meshes are rendered to produce binary images (silhouettes), where foreground pixels indicate visible parts of the shape.
-The error is then measured by taking the absolute value of the pixel-wise subtraction between the two silhouettes, averaged across all pixels and viewpoints (we use 400 viewpoints).
-This metric evaluates the visual similarity between the ground truth and the reconstructed mesh.
-Finally, the number of connected components measures global consistency and should be $1$ for our dataset of isolated plants.
-Figure 6: Qualitative results for 3D reconstruction by our method and SIREN [44], on synthetic clean, noisy and partial point clouds.
-We compare our method to SIREN [44], an implicit neural representation method that uses sine activation functions to model continuous signals, including signed distance functions (SDFs).
-For 3D reconstruction, the surface can be extracted as the zero level set of the SDF using Marching Cubes.
-SIREN was identified as the best-performing method for reconstructing plant geometry from point clouds by Prasad et al.
-[38], due to its ability to capture fine geometric details.
-Our method produces a complete smooth mesh that captures the plant visual shape and is robust with respect to input noise and missing input.
-Table 2 shows the results, where one can note that SIREN performs well in terms of the 3D distance-based error measures accuracy, completeness and Hausdorff distance.
-SIREN’s performance degrades significantly for depth images, and even more for noisy data.
-While our method has higher numerical error for these measurements than SIREN, it is much more robust to partial observations and noise.
-When considering silhouette loss, our method significantly outperforms SIREN as the visual similarity of the results is much higher for our method, (see Fig.
-6).
-While the number of connected components of the reconstruction is always $1$ for our method by construction, SIREN reconstructs highly fragmented shapes.
-Fig.
-6 shows examples of SIREN’s and our reconstructions for the three types of test examples.
-Note that while our reconstructions may have some misalignments, they are well structured and similar to the ground truth.
-In contrast, SIREN reconstructs fragmented shapes close to the input point clouds that are globally dissimilar from the ground truth plants.
-6.3 Skeletonization
-Synthetic Noise Depth images
-[53] $0.561\pm 0.312$ $0.701\pm 0.344$ $0.595\pm 0.324$
-[6] $0.544\pm 0.286$ $0.583\pm 0.301$ $0.716\pm 0.377$
-[30] $0.740\pm 0.355$ $0.776\pm 0.363$
-Ours $\mathbf{0.037\pm 0.052}$ $\mathbf{0.034\pm 0.044}$ $\mathbf{0.042\pm 0.055}$
-Table 3: Comparison to Xu et al. [53], Chaudhury et al. [6], and Livny et al. [30] for 3D skeleton extraction in terms of Chamfer Distance. “-” means that the method crashed due to numerical problems.
-We quantitatively evaluate our method for 3D skeleton extraction from input point cloud using the bidirectional Chamfer Distance.
-Chamfer Distance helps capturing the geometrical discrepancy between the ground truth skeleton and the predicted skeleton by measuring the sum of nearest neighbour distances in both directions.
-We compare our method with two classical plant skeletonization baselines Livny et al.[30], Xu et al.
-[53] and a stochastic skeleton refinement method Chaudhury et al.
-[6] that takes a predicted skeleton as input and outputs a refined skeleton.
-In this comparison we apply the skeleton refinement on the output skeletons from applying Xu et al.
-[53] on the plant point clouds.
-Figure 7: Qualitative results for skeletonization by our method, Livny et al.
-[30], Xu et al.
-[53], and Chaudhury et al.
-[6] refinement on Xu skeleton, from clean synthetic, noisy and partial point clouds taken from monocular depth maps.
-“-” means that the method crashed due to numerical problems.
-All these methods are designed to take scans of leafless plants as input.
-However, our dataset contains Chenopodium Album plants with leaves.
-Table 3 shows the results on the test sets of synthetic clean, noisy and partial point clouds.
-Our method shows consistent performance on the different test sets, suggesting robustness to noise and missing parts in the input point clouds for the skeletonization task.
-Since the baseline methods are not designed to take leaves as input, they tend to interpret leaf geometry as part of the skeleton, while the ground truth skeleton mainly describes the skeletal structure of the stem, branches and petioles.
-This explains the higher Chamfer distances, and the visual difference shown in Fig.
-7.
-6.4 Segmentation
-Stem Leaf Petiole Cotyledons
-Prec. Rec. F1 IoU Prec. Rec. F1 IoU Prec. Rec. F1 IoU Prec. Rec. F1 IoU
-Synthetic
-PlantNet 91 97 94 88 98 99 99 97 89 76 82 70 97 94 95 91
-PSegNet 91 99.5 95 90 99 99.6 99 98 95 88 92 85 99 89 94 88
-Ours 89 93 91 84 97 98 98 95 68 54 59 45 97 98 97 94
-Noise
-PlantNet 88 81 84 73 93 99 96 91 75 32 45 29 74 89 80 68
-PSegNet 99 75 85 74 90 99.7 95 90 76 21 33 20 98 80 88 79
-Ours 85 90 88 78 97 97 97 94 60 47 51 37 93 89 91 84
-Table 4: Comparison to PlantNet [26] and PSegNet [25] for semantic segmentation on synthetic and noisy points test sets using standard classification measures Precision, Recall, F1-Score and Intersection over Union (IoU).
-All values are percentages.
-We quantitatively evaluate our method’s performance on the semantic segmentation task applied to our test sets using standard classification metrics: precision, recall, F1 score, and Intersection over Union (IoU).
+10 Additional evaluations
+10.1 Semantic Segmentation
+We quantitatively evaluate our method’s performance on the semantic segmentation task applied to our test sets using standard classification metrics precision, recall, F1 score, and Intersection over Union (IoU).
 For each class, precision is defined as the ratio of true positives to the sum of true positives and false positives, while recall is the ratio of true positives to the sum of true positives and false negatives.
@@
 Finally, IoU is computed as the ratio of the intersection to the union of the predicted and ground truth point sets for a given class.
-We compare our method to the two strong baselines PlantNet [26] and PSegNet [25].
-Both PlantNet and PSegNet take a deep neural network approach with special sampling strategies, and are suitable for semantic and instance segmentation of plants.
-Table 4 shows the results on the test sets of synthetic and noisy point clouds.
-We do not test segmentation methods on the depth image test set, as we do not have ground truth labels in this case.
-For synthetic scans, PSegNet outperforms all the methods except on Cotyledons segmentation, where our method shows a better performance.
-However, when noise is introduced, our method outperforms the other methods for most of the measures.
-Note that the petiole is the most difficult part to segment; this is because of its cylindrical shape, which is very close to the stem branches and makes it difficult to discriminate between these two classes.
-Overall, our method is robust to noise and performs on-par with the strong baselines.
+Table 4 shows the results on all test sets.
+Our method is robust to noise and performs overall on-par with the strong baselines PlantNet [LI2022243] and PSegNet [psegnet2022] on each individual task, while performing all the tasks at once.
+Stem Leaf Petiole Cotyledons
+Prec. Rec. F1 IoU Prec. Rec. F1 IoU Prec. Rec. F1 IoU Prec. Rec. F1 IoU
+Clean
+PlantNet 85.3 88.2 86.7 76.6 98.4 99.5 98.9 97.9 82.1 71.7 76.5 62.0 98.9 94.4 96.7 93.5
+PSegNet 99.1 99.3 99.2 98.5 98.4 98.9 98.6 97.3 66.6 59.6 62.9 45.9 99.7 97.0 98.6 97.3
+Ours 84.8 88.0 86.2 75.9 98.0 98.4 98.2 96.5 66.7 59.4 62.2 45.9 96.7 95.8 96.2 92.7
+Noisy
+PlantNet 84.2 88.4 86.2 75.8 97.7 99.1 98.4 96.9 77.6 62 69.0 52.6 96.4 93.1 94.7 89.9
+PSegNet 99.3 83.6 90.7 83.1 95.9 99.9 97.9 95.9 96.2 29.6 45.3 29.3 99.6 98.1 98.9 97.7
+Ours 83.0 86.3 84.5 73.3 97.9 98.3 98.1 96.2 64.2 57.4 59.9 43.7 95.8 93.9 94.8 90.2
+Depth maps
+PlantNet 86.2 86.8 86.5 76.2 98.5 99.2 98.8 97.7 70.2 65.6 67.8 51.3 97.4 88.7 92.9 86.7
+PSegNet 99.1 98.9 99.0 98.1 96.2 99.9 98.1 96.2 98.9 54.9 70.6 54.6 99.8 98.7 99.2 98.5
+Ours 84.6 84.2 84.0 73.2 98.4 98.6 98.5 97.0 57.9 54.1 54.5 38.8 96.8 94.8 95.7 92.0
+Table 4: Comparison to PlantNet [LI2022243] and PSegNet [psegnet2022] for semantic segmentation on clean and noisy points test sets using standard classification measures Precision, Recall, F1-Score and Intersection over Union (IoU).
+All values are percentages.
+10.2 Instance Segmentation
+We qualitatively compare our method to PlantNet [LI2022243] and PSegNet [psegnet2022] for the Instance segmentation task.
 Fig.
-8 shows the semantic segmentation outputs of our method compared to the ground truth, where the semantics are properly captured even for noisy inputs.
-Furthermore, we show our method’s ability to perform instance segmentation in Fig.
-9, where it successfully separates individual plant parts and remains robust to noise.
-Figure 8: Examples of semantic segmentation results on synthetic clean and noisy point clouds.
-Figure 9: Examples of instance segmentation results on synthetic clean and noisy point clouds.
-7 Conclusion and Discussion
-In this paper we have presented a data-driven method to infer plants parametric representations i.e.
-L-Strings, from 3D unstructured point cloud input.
-We have shown that the L-String representation contains structural and geometric information that the input scans lack, and that such information allows for multiple tasks like 3D reconstruction, skeletonization and segmentation.
-Our results have shown that our method is on-par with the state-of-the-art for these tasks on synthetic data, and is robust to noisy and incomplete input.
-However, the performance of our method is limited to synthetic data, and first results from this first version of the method need to be qualitatively and quantitatively improved when applied to real scans.
-Figure 10: Our 3D reconstruction results on input scans of real Chenopodium Album plants.
-This work demonstrate the feasibility to learn latent space of 3D plant structures and reconstruct plant architecture parametric models from it.
-In the next step, our aim is to upscale our method to address fully real 3D plant point clouds.
-For this we need to investigate the level of accuracy required to simulated plants real plants so that the method can use simulated plants to learn real plant architectures.
-Fig 10 shows first tests of our 3D reconstructions from real scans.
-Although the reconstructions are not yet accurate on all levels, global information is learned by our model.
-8 Acknowledgements
-We thank Franck Hétroy-Wheeler for helpful discussions, and Ayan Chaudhury, Frédéric Boudon, and Aymen Merrouche for providing support for their code.
-This work was partially supported by French government funding managed by the National Research Agency under grant ANR-24-CE23-1586 (4DPlants).
+9 shows the instance segmentation results on examples from all test sets, where each individual plant organ is assigned a unique color.
+Note that the colors are to distinguish the segmented organs without any correspondence between the results and the ground truth.
+Our method is on-par with the baselines for this task where it is able to segment individual organ instances and is robust to noise and partial data.
+Figure 9: Visual comparison between our method, PlantNet and PSegNet for the instance segmentation task.
+Different colors are assigned to segmented instances without correspondence to the ground truth.
+10.3 Skeletonization
+To support the results shown in Table 3 in the main paper on the skeletonization task, we qualitatively compare to the baselines Livny et al.
+[livny2010automatic], Xu et al.
+[10.1145/1289603.1289610] and Chaudhury et al.
+[Chaudhury2020.02.15.950519] on all test sets.
+Fig.
+11 shows the visual qualitative results of all methods on extracting a full skeleton from an input point cloud, with our results of extracting the branch skeleton compared to the ground truth.
+Our results are close to the ground truth skeletons and are robust to noise and partial input on both scales.
+10.4 Influence of Latent Dimension
+All loss functions are necessary by construction for our models to train and thus cannot be ablated.
+Instead, we study the influence of the dimension of the learned latent space $\mathcal{S}$ on the performance of our model.
+For that, we evaluate the L-string latent representation stage by training the recursive auto-encoders with different latent dimensions $dim_{\mathcal{S}}$ and measuring the Tree-Edit Distance [ferraro-hal-00827474] between the input L-strings and the reconstructions.
+Tree-Edit distance measures the difference between two tree graphs w.r.t.
+the number of operations needed to translate from one tree structure to the other, with additional local cost functions that is the difference between the node parameters in our case, hence it measures the difference in both topology and shape between plants represented as L-strings.
+Table 5 shows the average reconstruction error in tree-edit distance for different dimensions, in our implementation we choose $dim_{S}$ that is most compact with minimum error.
+$dim_{\mathcal{S}}$ Tree Edit Distance
+32 0.227
+64 0.196
+128 0.197
+256 0.296
+Table 5: Analysis of the influence of latent dimension on the performance of the L-string auto-encoder model w.r.t. the Tree-Edit Distance.
+10.5 Latent Space Analysis
+To analyse the data distribution in the learned latent space $\mathcal{S}$, we project the latent points representing plants from the training set in Fig.
+10 using UMAP [mcinnes2020umapuniformmanifoldapproximation], the uniform manifold approximation and projection non-linear technique for dimension reduction with local and global structure preservation.
+Different plant structures are assigned different colors in Fig.
+10, one can notice that our model learns to encode plants sharing the same tree structure close by in $\mathcal{S}$.
+Figure 10: Analysis of the distribution of different tree structures from the training data in the latent space projected in 2D using UMAP.
+Each color represent a unique plant structure in the training data.
+Plants sharing the same structure form clusters in $\mathcal{S}$.
+Figure 11: Comparison to Livny et al[livny2010automatic], Xu et al[10.1145/1289603.1289610], and Chaudhury et al[Chaudhury2020.02.15.950519] refinement on Xu skeleton for skeletonization.
+“-” means that the method crashed due to numerical problems.
+Note that ours is the only method applicable at different scales, and able to output a full skeleton or a branch skeleton, while achieving visually accurate results.